### v2

# Collusion-Resilience in Transaction Fee Mechanism Design

Abstract
Users bid in a transaction fee mechanism (TFM) to get their transactions included and confirmed by a blockchain protocol. Roughgarden (EC’21) initiated the formal treatment of TFMs and proposed three requirements: user incentive compatibility (UIC), miner incentive compatibility (MIC), and a form of collusion-resilience called OCA-proofness. Ethereum’s EIP-1559 mechanism satisfies all three properties simultaneously when there is no contention between transactions, but loses the UIC property when there are too many eligible transactions to fit in a single block. Chung and Shi (SODA’23) considered an alternative notion of collusion-resilience, called $c$-side-contract-proofness ($c$-SCP), and showed that, when there is contention between transactions, no TFM can satisfy UIC, MIC, and $c$-SCP for any $c\geq 1$. OCA-proofness asserts that the users and a miner should not be able to “steal from the protocol.” On the other hand, the $c$-SCP condition requires that a coalition of a miner and a subset of users should not be able to profit through strategic deviations (whether at the expense of the protocol or of the users outside the coalition).
Our main result is the first proof that, when there is contention between transactions, no (possibly randomized) TFM in which users are expected to bid truthfully satisfies UIC, MIC, and OCA-proofness. This result resolves the main open question in Roughgarden (EC’21). We also suggest several relaxations of the basic model that allow our impossibility result to be circumvented.
1 Introduction
Real estate on the blockchain is scarce, and blockchain users bid in an auction called the transaction fee mechanism (TFM) to have their transactions included and confirmed on the blockchain. The original Bitcoin protocol adopted a simple first-price auction, where the top $k$ bids win and they each pay their bid. However, such first-price auctions are known to incentivize untruthful bidding. Therefore, a line of subsequent works [LSZ19, Yao, BEOS19, BCD+, Rou20, Rou21, FMPS21, CS23, SCW23, WSC24, GY22, ZCZ22, BGR23, TY23, KKLP23, XFP23, CMW23, LRMP23, Ndi23] explored what is the “dream TFM” for blockchains. Most works [Rou20, Rou21, CS23, SCW23, WSC24, GY22, GY22, ZCZ22, BGR23, TY23] agree on roughly the same set of desiderata, that is, a dream TFM should provide incentive compatibility not just for an individual user, but also for the miner of the block. Further, a dream TFM should provide resilience against miner-user collusion.
Roughgarden [Rou21] was the first to formally define the aforementioned requirements for a TFM, which he referred to as user incentive compatibility [^1] User incentive compatibility (UIC) is usually called dominant-strategy incentive compatible (DSIC) in the mechanism design literature. In general, we allow UIC TFMs to make use of non-truthful (but dominant) bidding strategies (see Definition 2)., (myopic) miner incentive compatibility, and OCA-proofness, where OCA stands for “off-chain agreement” and refers to colluding strategies between the miner and a set of users that allow offchain transfers. Roughgarden [Rou21] also showed that the simple “posted price auction with all fees burnt” mechanism, which corresponds to the behavior of Ethereum’s EIP-1559 TFM [BCD+] when there is no congestion, satisfies all three properties. However, the posted price auction with all fees burnt does not satisfy all three properties when there is congestion. In practice, congestion does occur when there are major events such as an NFT mint or price fluctuations — for example, in Ethereum, roughly 2.3% of the blocks experience congestion. [^2] From Jan 1, 2024 to Feb 5, 2024, 256595 blocks have been produced on Ethereum, and 5840 blocks among them were full (meaning more than 99.9% of the gas limit (30M) was used). When congestion arises, approximately speaking, Ethereum’s EIP-1559 mechanism falls back to the first-price auction, violating user incentive compatibility. Therefore, an interesting question is whether we can design a dream TFM satisfying all three properties for finite block sizes.
Chung and Shi [CS23] considered an alternative notion of collusion-resilience, called side-contract-proofness. Unfortunately, they proved that no (even randomized) TFM can simultaneously satisfy user incentive compatibility and side-contract-proofness. Because side-contract-proofness is a more demanding property than OCA-proofness, the question raised by Roughgarden [Rou21], of whether there is a dream TFM satisfying all three properties under his collusion-resilience notion, had remained open.
Two notions of miner-user collusion-resilience.
Multiple natural notions of collusion-resilience can and have been studied in the context of TFM design. Here we clarify informally the key differences between the notions proposed by Roughgarden [Rou21] and Chung and Shi [CS23]. These notions are defined formally in Definitions 46 (see Section 2) and compared further via examples in Appendix A.
  • OCA-proofness: Roughgarden’s notion, henceforth referred to as OCA-proofness, asserts that there should exist a “reference strategy” for a miner and all users that is guaranteed to maximize their joint surplus. In this reference strategy, the miner is expected to follow the inclusion rule intended by the TFM. For users, the definition requires only that users follow some fixed bidding strategy $\sigma$ (i.e., a mapping from a private user valuation to a user bid) that is individually rational (i.e., $\sigma(v)\leq v$ for all $v\geq 0$). In particular, in the reference strategy, users are expected to bid independently (with a user’s bid independent of other users’ valuations and bids), and expected to submit a single bid (with no additional fake bids injected). One example of such a bidding strategy is the truth-telling strategy (with $\sigma(v)=v$). Because Roughgarden [Rou21] wished to discuss the OCA-proofness properties of non-UIC TFMs like first-price auctions, the definition also allows the reference strategy to be defined by a non-truthful bidding strategy (e.g., $\sigma(v)=v/2$). As a consequence, to prove that a TFM is both UIC and OCA-proof, it is sufficient to prove that it is UIC under one bidding strategy and OCA-proof under a possibly different bidding strategy (as in the example in Section 6.1).
  • $c$-SCP: Chung and Shi’s notion [CS23], henceforth called $c$-SCP (where SCP stands for side-contract-proofness), requires that the honest strategy (i.e., all users follow the honest bidding rule and the miner honestly implements the inclusion rule) is the profit-maximizing strategy for any coalition consisting of the miner of the present block and at most $c$ users. For truthful mechanisms, the honest bidding rule is the truthful one, while for non-truthful mechanisms, the bidding rule can be more general (see Section 2.1 for the formal definition). Chung and Shi’s notion aligns with standard notions used in a line of work at the intersection of game theory and cryptography  [HT04, KN08, ADGH06, OPRV09, AL11, ACH11, GKM+13, GKTZ15, GTZ15, Kat08, DR07, GLR10, CGL+18, WAS22, CCWS21, PS17, KMSW22, FW20, EFW22].
Discussion.
The two notions of collusion-resilience address different issues. OCA-proofness captures the intuitive requirement that the users and miners should not be able to steal from the protocol through strategic deviations — for this reason, it considers only the global coalition consisting of the miner and all users. By contrast, the $c$-SCP notion captures the intuitive idea that a miner-user coalition’s best response is to act honestly, and that no strategic deviations can allow the coalition to steal from other users or steal from the protocol. For further discussion, see the end of this section and Appendix A.
1.1 Our Contributions
As explained above, both Roughgarden’s and Chung and Shi’s collusion-resilience notions capture meaningful incentive compatibility considerations. Recognizing their differences, one natural question is: does Chung and Shi’s finite-block impossibility result still hold if we adopt the original OCA-proofness notion of Roughgarden in lieu of $c$-SCP? Notably, no existing TFM construction [LSZ19, Yao, BEOS19, BCD+, Rou20, Rou21, FMPS21, CS23, SCW23, WSC24, GY22, ZCZ22, BGR23, TY23, KKLP23, XFP23, CMW23, LRMP23, Ndi23] simultaneously satisfies user incentive compatibility, miner incentive compatibility, and OCA-proofness under finite block size.
Main impossibility result.
In our work, we give an affirmative answer to the above question. We show that, indeed, an analog of Chung and Shi’s finite-block impossibility result still holds when we replace the $c$-SCP requirement with OCA-proofness. Specifically, we prove the following theorem.
Theorem 1.1.
Suppose the block size is finite. Then, no possibly randomized, truthful TFM can simultaneously satisfy user incentive compatibility (UIC), miner incentive compatibility (MIC), and OCA-proofness. Further, this impossibility holds even when the globally optimal strategy $\sigma$ need not be individually rational.
In a truthful TFM, a user is expected to bid truthfully, so if the mechanism satisfies UIC, a user’s utility is maximized when it just reports its true value. However, OCA-proofness allows the global coalition to adopt a non-truthful bidding strategy $\sigma$ even for truthful mechanisms.
Our Theorem 1.1 is intuitively stronger but technically incomparable in comparison with Chung and Shi’s impossibility, which shows that no TFM can simultaneously satisfy UIC and 1-SCP for finite block sizes. The reason is that Chung and Shi’s impossibility does not rely on MIC; however, MIC is necessary for our Theorem 1.1 to hold. Specifically, a simple second-price auction with no burning (see Remark 2) satisfies both UIC and OCA-proofness, but does not satisfy MIC since the miner may benefit by injecting a fake $(t+1)$-th bid where $t$ is the number of confirmed bids, since the $(t+1)$-th bid sets the price for confirmed bids.
Global SCP.
We suggest a simpler version of OCA-proofness that we call global SCP, which also intuitively captures the requirement that strategic users and miners cannot steal from the protocol, and is perhaps more appropriate when focusing on UIC TFMs (as we do in this paper). In our work, global SCP is not only a technical stepping stone towards proving Theorem 1.1, but also of independent interest as we explain below. Specifically, global SCP is almost the same as OCA-proofness, except for requiring $\sigma$ to be the honest bidding strategy indicated by the mechanism (i.e., the same bidding strategy used to establish UIC). In other words, a mechanism satisfies global SCP if and only if the honest strategy is surplus-maximizing for the global coalition. It is easy to see that for a truthful mechanism, $c$-SCP for any $c$ implies global SCP, which in turn implies OCA-proofness. To prove Theorem 1.1, we first prove the following theorem:
Theorem 1.2.
Suppose that the block size is finite. Then no possibly randomized TFM can simultaneously satisfy user incentive compatibility (UIC), miner incentive compatibility (MIC), and global SCP. Further, the impossibility holds even for non-truthful mechanisms.
We now explain why the global SCP notion is of independent interest. One advantage of global SCP is that the revelation principle holds for any TFM that satisfies UIC, MIC, and global SCP, which we formally prove in Section 8. In other words, given any TFM that is UIC, MIC, and global SCP, there is an equivalent truthful mechanism that simulates it. For this reason, Theorem 1.2 rules out even non-truthful TFMs that simultaneously satisfy UIC, MIC, and global SCP. [^3] Simultaneously with and independently of this paper, Gafni and Yaish [GY24] proved, among other results, a version of Theorem 1.2 for the special case of deterministic mechanisms and a block size of 1.
By contrast, Theorem 1.1 holds only for truthful mechanisms. In particular, in Section 6.1, we show a non-truthful mechanism that simultaneously satisfies UIC, MIC, and OCA-proof. The mechanism is contrived, but it demonstrates the subtlety and the technical challenges when modeling the notion of collusion-resilience. This also suggests that the revelation principle does not hold for mechanisms that satisfy UIC, MIC, and OCA-proofness, partly because in such a mechanism, the bidding strategies used to establish UIC and OCA-proofness may be different.
Ways to circumvent the impossibilities.
We show in Section 7 that the impossibility of Theorem 1.1 can be circumvented by allowing non-truthful mechanisms or by allowing users to coordinate in bidding in the globally optimal strategy $\sigma$. In the same section, we raise an open question regarding whether it is possible to use cryptography (e.g., the MPC-assisted model of Shi et al. [SCW23]) and Bayesian notions of incentive compatibility to circumvent the impossibilities.
1.2 TFM Incentive-Compatibility Notions: A Cheat Sheet
We gather here informal definitions and comparisons of the key incentive-compatibility notions used in this paper. First, a transaction fee mechanism specifies how a user is supposed to bid (as a function of its private valuation), which transactions a miner is supposed to include (as a function of the transactions it knows about and their bids), and the resulting outcome (the subset of included transactions that get confirmed, and the payments made by the users and received by the miner). If the bidding strategy suggested by the TFM is the identity, then we additionally call the TFM truthful. In this paper, as in the rest of the TFM literature, we consider only static mechanisms.
  • UIC. (Definition 2) Provided that the miner follows the suggested inclusion rule, the bidding strategy suggested by the TFM is a dominant strategy for users.
  • MIC. (Definition 3) The inclusion rule suggested by the TFM is always revenue-maximizing for the miner regardless of users’ bids; moreover, the miner cannot increase its revenue through the injection of fake transactions.
  • Global SCP. (Definition 5) If the miner follows the inclusion rule suggested by the TFM and all users follow the bidding rule suggested by the TFM, then their joint surplus is at least as large as it would be from any coordinated deviation.
  • $c$-SCP. (Definition 4) For every coalition of the miner and at most $c$ users, if the miner follows the inclusion rule suggested by the TFM and the users in the coalition follow the bidding rule suggested by the TFM, then the joint surplus of the coalition is at least as large as it would be from any coordinated deviation (holding fixed the bids of users outside the coalition).
  • OCA-proof. (Definition 6) If the miner follows the inclusion rule suggested by the TFM and all users follow a suitably chosen individually rational bidding rule $\sigma$ (possibly different from the one suggested in the TFM description), then their joint surplus is as large as it would be from any coordinated deviation.
For example, in [Rou21] it was shown that Ethereum’s EIP-1559 TFM and a variant called the “tipless mechanism” satisfy UIC, MIC, and OCA-proofness when there is no contention between transactions; in fact, in this case, these TFMs satisfy the $c$-SCP condition for every $c\geq 1$. When there is contention between transactions for inclusion in a block, the EIP-1559 TFM loses its UIC property and the tipless mechanism loses (all three notions of) collusion-resilience.
As mentioned above:
  • (Theorem 8.1) A relevation principle holds for the global SCP and $c$-SCP notions: any UIC and MIC TFM that satisfies one of these properties can be simulated by a truthful UIC and MIC TFM that satisfies the same property.
  • A relevation principle does not in general hold for the OCA-proof notion: while there are non-truthful TFMs that satisfy UIC, MIC, and OCA-proofness (Section 6.1), there are no truthful TFMs that satisfy UIC, MIC, and OCA-proofness (Theorem 6.9).
The main result in Chung and Shi [CS23] states that, even among randomized TFMs, no TFM satisfies UIC and $c$-SCP for any $c\geq 1$. Our Theorem 5.6 proves that, even among randomized TFMs, no TFM satisfies UIC, MIC, and global SCP. (Due to the revelation principle mentioned above, these impossibility results apply to both truthful and non-truthful TFMs.) Our Theorem 6.9 proves the stronger statement that, even among randomized TFMs, no truthful TFM satisfies UIC, MIC, and OCA-proofness.
Reflecting on the competing notions of collusion-resilience, we can observe the following. The $c$-SCP condition may be particularly appropriate in scenarios where the primary concern is deviations by small coalitions, or in scenarios where users may wish to deviate in ways that exploit other users. The $c$-SCP condition is also notable in that, together with the UIC condition, it already triggers the impossibility result in [CS23] (without any appeal to MIC). The OCA-proofness condition is distinguished by being the weakest of the three notions (thus leading to the strongest impossibility results) and by allowing the discussion of non-UIC mechanisms. [^4] For example, in a first-price auction, the “reference outcome” might be defined by a (non-truthful) bidding strategy that would constitute a Bayes-Nash equilibrium with respect to some prior over user valuations (cf., Corollary 5.12 in [Rou21]). For TFMs that are UIC and MIC, like those studied in this paper, global SCP is arguably the “right” definition—capturing the spirit of OCA-proofness, without any additional technical complications arising from users using different bidding strategies to satisfy UIC and collusion-resilience. Put differently, the UIC and MIC conditions imply that the miner and the users following their intended strategies constitutes a Nash equilibrium; the global SCP condition asserts that this Nash equilibrium is also robust to deviations by the grand coalition, while OCA-proofness only asserts such robustness for a possibly different strategy profile (defined by the intended inclusion rule and some individually rational bidding strategy). From this vantage point, one might view Theorem 5.6 as the main impossibility result in this paper, with Theorem 6.9 serving as a technically challenging extension of the result under still weaker incentive-compatibility conditions.
2 Definitions
2.1 Transaction Fee Mechanism
In this work, we assume each user has a true value $v\in\mathbb{R}_{\geq 0}$ for its transaction being confirmed on the blockchain, where $\mathbb{R}_{\geq 0}$ denotes non-negative reals. We assume all transactions have an equal size, and we define the block size to be the maximal number of transactions that can be included in a block. If there exists a finite number $k$ such that a block can include at most $k$ transactions, we say the block size is finite; [^5] The finite block size regime in this work and [CS23] corresponds to the case in [Rou21] where the base fee in the EIP-1559 or tipless mechanisms is excessively low, i.e. the number of transactions willing to pay the base fee exceeds the maximum block size (cf., Definition 5.6 in [Rou21]). otherwise, we say the block size is infinite. When the block size is finite, we always assume $k\geq 1$. Additionally, we assume that the users’ utilities are independent of the order of the transactions in a block. In practice, transactions may have different sizes and the order of the transactions matters. For example, the size is measured in “gas” in Ethereum, and the prices for trading may depend on the order of the transactions. However, for our impossibility results, these assumptions only make them stronger.
A transaction fee mechanism (TFM) consists of the following possibly randomized algorithms:
  • Bidding rule (executed by the user): takes in the user’s true valuation $v\in\mathbb{R}_{\geq 0}$ for getting its transaction confirmed, and outputs a vector of a non-negative number of real-valued bids. Without loss of generality, we assume that at most one bid in the output vector may correspond to the user’s actual transaction [^6] The blockchain protocol can always suppress conflicting or double-spending transactions. which has a true value of $v$ (henceforth called the primary bid), and the remaining bids are fake bids with a true value of 00. We say a mechanism is truthful if the honest bidding rule asks the user to submit a single primary bid that reports its true value. Otherwise, if the honest bidding rule may ask the user to submit bids other than its true value, the mechanism is said to be non-truthful. Our formulation is more general and admits non-truthful mechanisms, too. [^7] Throughout the paper except Section 8, we only focus on bidding rules that output a single bid. In Section 8, we consider general bidding rules that may output multiple bids.
  • Inclusion rule (executed by the miner): takes a bid vector ${\mathbf{b}}=(b_{1},\dots,b_{t})\in\mathbb{R}^{t}$ as input, and outputs a subset of $S\subseteq{\mathbf{b}}$ indicating which bids to be included in the block. When the block size $k$ is finite, it must be that $|S|\leq k$.
  • Confirmation rule (executed by the blockchain): takes as input a block $\mathbf{B}$, i.e., a vector of included bids, and outputs a subset of $\mathbf{B}$ to be confirmed. In general, not all transactions included in the block must be confirmed, and only the confirmed transactions are executed. [^8] Roughgarden [Rou21] assumes that all included transactions are confirmed. However, Chung and Shi [CS23] show that allowing unconfirmed transactions in a block enlarges the design space. For example, some mechanisms require a block to contain some unconfirmed transactions (see Section 7 in [CS23]).
  • Payment rule (executed by the blockchain): given a block $\mathbf{B}$ as input, outputs the payment of each bid. We require individual rationality, i.e., for any vector of true values $\mathbf{v}=(v_{1},\ldots,v_{t})$, if all users execute the honest bidding rule, and we then execute the honest inclusion, confirmation, and payment rules, then the following holds with probability $1$: [^9] We can also relax the requirement such that individual rationality holds in expectation. Both the impossibility results (Sections 4, 5 and 6.2) and the revelation principle result (Section 8) continue to hold. for every user $i\in[t]$, if its primary bid is not confirmed, then its total payment is zero; else if its primary bid is confirmed, then its total payment is at most its true value $v_{i}$. Here, a user’s total payment is the sum of the payment of all bids it has submitted.
  • Miner revenue rule (executed by the blockchain): given a block $\mathbf{B}$ as input, outputs how much revenue the miner gets. We require budget feasibility, i.e., the miner’s revenue is at most the total payment collected from all confirmed bids. When the miner revenue is strictly smaller than the total payment, the difference between the total payment and the miner revenue is said to be burnt, i.e. eliminated from circulation.
We say a TFM is trivial if the confirmation probability of all transactions is zero for any bid vector assuming the miner honestly follows the inclusion rule; otherwise, it is called non-trivial.
A strategic miner or miner-user coalition may deviate from the honest inclusion rule. On the other hand, since the confirmation, payment, and miner revenue rules are executed by the blockchain, they are always implemented honestly.
We focus on mechanisms that are weakly symmetric, i.e., mechanisms that do not make use of the bidders’ identities or other auxiliary information (e.g., timestamp, transaction metadata), except for tie-breaking among equal bids. More formally, we define weak symmetry as below.
Definition 1 (Weak symmetry).
A mechanism is called weakly symmetric if the mechanism can always be equivalently described in the following manner: given a bid vector ${\mathbf{b}}$ where each bid may carry some extra information such as identity or timestamp, the honest mechanism always sorts the vector ${\mathbf{b}}$ by the bid amount first. During the sorting step, if multiple bids have the same amount, then arbitrary tie-breaking rules may be applied, and the tie-breaking can depend on extra information such as timestamp, identity, or random coins. After this sorting step, the inclusion rule and the confirmation rules should depend only on the amount of the bids and their relative position in the sorted bid vector.
The above weak symmetry definition is equivalent to the following: for two bid vectors ${\mathbf{b}}$ and ${\mathbf{b}}^{\prime}$ of length $n$ that are equivalent up to a permutation, then the distribution of the mechanism’s outcomes on ${\mathbf{b}}$ and ${\mathbf{b}}^{\prime}$ are identically distributed — assuming that the outcome is encoded in the form of $\left(\mu,{\sf sorted}(\{(v_{i},x_{i},p_{i})\}_{i\in[n]})\right)$ sorted in descending order of $v_{i}$, where $\mu$ encodes the miner’s revenue, each tuple $(v_{i},x_{i},p_{i})$ indicates that some user with bid $v_{i}$ had the confirmation outcome $x_{i}\in\{0,1\}$, and paid $p_{i}$.
If we require that the tie-breaking cannot depend on the extra information, we can define strong symmetry as follows: for any bid vector ${\mathbf{b}}=(b_{1},\dots,b_{t})$, and for any $i\neq j$ such that $b_{i}=b_{j}$, the random variables $(x_{i},p_{i})$ and $(x_{j},p_{j})$ must be identically distributed. Throughout the paper, we focus on weakly symmetric mechanisms. Since a strongly symmetric mechanism must also be weakly symmetric, our impossibility results also rule out the existence of strongly symmetric mechanisms.
Strategy space.
A strategic user can deviate from the honest bidding rule and post an arbitrary bid vector with zero to multiple bids. Without loss of generality, we may assume that in the strategic bid vector, at most one bid can correspond to the user’s actual transaction which has a non-zero true value; all other bids must be fake bids with zero true value. A strategic miner can deviate from the honest inclusion rule, and instead create an arbitrary block (subject to the block size limit) that includes any subset of the bid vector as well as any number of fake bids that it chooses to inject. A strategic miner-user coalition can adopt a combination of the above strategies.
Utility and social welfare.
For a user with true value $v$, let $x\in\{0,1\}$ be the indicator of whether its primary bid is confirmed or not, let $p$ denote its total payment, then the user’s utility is $x\cdot v-p$. The miner’s utility is simply its revenue. The social welfare is defined to be the sum of the utilities of all users and the miner (i.e., the total value of the confirmed transactions, less any burned payments).
Notice that we allow the miner revenue to be smaller than the sum of users’ payment, since the coins can be burnt. When calculating the social welfare, the payments among the users and the miner are canceled out, so the social welfare is independent of the payment; however, the amount of burnt coins decreases the social welfare. For example, suppose there is only one user, and let $p$ be the user’s payment and $q$ be the amount of burnt coins. In this case, the user’s utility is $x\cdot v-p$, the miner revenue is $p-q$, and the social welfare is $(x\cdot v-p)+(p-q)=x\cdot v-q$.
2.2 Incentive Compatibility Notions
Definition 2 (User incentive compatible (UIC)).
A TFM is said to be user incentive compatible (UIC), iff the following holds: for any user $i$, for any bid vector ${\mathbf{b}}_{-i}$ that corresponds to the bids posted by all other users, user $i$’s expected utility is always maximized when it follows the honest bidding rule (assuming that the miner executes the inclusion rule honestly). Notice that for a truthful mechanism, the bidding rule is just truth-telling, so following the honest bidding rule simply means submitting the value truthfully; for a non-truthful mechanism, a bidding rule may ask a user to submit a bid other than its true value.
Definition 3 (Miner incentive compatible (MIC)).
A TFM is said to be miner incentive compatible (MIC), iff given any bid vector ${\mathbf{b}}$, the miner’s expected utility is maximized when the miner does not inject any fake bid and creates a block indicated by the honest inclusion rule.
Definition 4 ($c$-side-contract-proof ($c$-SCP)).
A TFM is said to be $c$-side-contract-proof ($c$-SCP), iff for any coalition $C$ consisting of the miner and between $1$ and at most $c$ users, and for any bid vector ${\mathbf{b}}_{-C}$ that corresponds to the bids posted by all users not in $C$, the coalition’s joint utility is always maximized when it adopts the honest strategy, i.e., all users in the coalition bid according to the honest bidding rule, and the miner follows the honest inclusion rule.
Definition 5 (Global side-contract-proof (global SCP)).
A TFM is said to be global side-contract-proof (global SCP), iff given any vector of true values $\mathbf{v}$, the expected social welfare is maximized when all the users bid according to the honest bidding rule, and the miner follows the honest inclusion rule, where the maximization is taken over all the coordinated strategies that the coalition consisting of the miner and all users can adopt.
Definition 6 (OCA-proof).
A TFM is said to be OCA-proof iff there exists an individually rational bidding strategy $\sigma:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}$ such that the expected social welfare is maximized when all the users bid according to $\sigma$, and the miner creates a block indicated by the inclusion rule, where the maximization is taken over all the coordinated strategies that the coalition consisting of the miner and all users can adopt.
In the definitions above, the expectation is taken over the randomness of the TFM. More explicitly, in Definition 2, the expectation is taken over the randomness of the inclusion/confirmation/payment rules; in Definitions 3, 4, 5 and 6, the expectation is taken over the randomness of the inclusion/confirmation/
payment/miner revenue rules.
Note that in the OCA-proofness definition, $\sigma$ is required to output a single real-valued bid. A canonical example of $\sigma$ is scaling; that is, $\sigma(v)=\gamma v$ for some $\gamma\in[0,1]$ (cf., Corollary 5.12 and 5.14 in [Rou21]).
A detailed comparison between $c$-SCP, global SCP, and OCA-proofness is given in Appendix A.
3 Preliminary: Myerson’s Lemma
If a TFM satisfies UIC, then for any user, its confirmation probability and the payment must satisfy the famous Myerson’s lemma [Mye81]. It is an important ingredient for proving our impossibilities, and we review it in this section. Formally, given a bid vector ${\mathbf{b}}=(b_{1},\dots,b_{t})$, let $x_{i}({\mathbf{b}})\in[0,1]$ denote the probability that user $i$’s bid is confirmed, and let $p_{i}({\mathbf{b}})$ denote user $i$’s expected payment assuming the miner implements the mechanism honestly. For any vector ${\mathbf{b}}$, let ${\mathbf{b}}_{-i}$ denote a vector when we remove $i$-th coordinate in ${\mathbf{b}}$. With these notations, Myerson’s lemma can be stated as follows.
Lemma 3.1 (Myerson’s lemma [Mye81]).
For any TFM that satisfies UIC:
  • For any user $i$, bid vector ${\mathbf{b}}_{-i}$, and $b_{i},b^{\prime}_{i}$ such that $b_{i}^{\prime}>b_{i}$, it must be $x_{i}({\mathbf{b}}_{-i},b^{\prime}_{i})\geq x_{i}({\mathbf{b}}_{-i},b_{i})$.
  • For any user $i$, bid vector ${\mathbf{b}}_{-i}$ from other users, and bid $b_{i}$ from user $i$, user $i$’s expected payment must be
[…]
    with respect to the normalization condition: $p_{i}({\mathbf{b}}_{-i},0)=0$, i.e. user $i$’s payment must be zero when $b_{i}=0$.
When the mechanism is deterministic, the confirmation probability $x_{i}$ is either 00 or $1$. In this case, user $i$’s payment can be simplified as
$p_{i}({\mathbf{b}}_{-i},b_{i})=\left\{\begin{matrix}\inf\{z\in[0,b_{i}]:x_{i}({\mathbf{b}}_{-i},z)=1\},&\text{ if $x_{i}({\mathbf{b}}_{-i},b_{i})=1$};\\ 0,&\text{ if $x_{i}({\mathbf{b}}_{-i},b_{i})=0$.}\end{matrix}\right.$
Conceptually, user $i$ must pay the minimal price which makes its bid confirmed.
4 Warmup: Impossibility of UIC + MIC + Global SCP for Deterministic Mechanisms
As a warmup, we first show a finite-block impossibility for UIC + MIC + global SCP for deterministic mechanisms. Recall that a TFM is said to be trivial if everyone’s confirmation probability is zero for any bid vector assuming the miner follows the inclusion rule. In this case, everyone’s utility is always zero in an honest execution. We will show that no non-trivial mechanism can satisfy all three properties simultaneously. Later in Section 5, we extend the impossibility to randomized mechanisms. Due to the revelation principle that we prove in Section 8, if we can prove the impossibility for truthful mechanisms, the impossibility immediately extends to non-truthful mechanisms as well. Therefore, in this section, we shall assume truthful mechanisms.
Lemma 4.1.
For any global SCP mechanism, the confirmed bids must correspond to the highest bids.
Proof.
Suppose in some scenario, Alice bids her true value $b$ and Bob bids his true value $b^{\prime}<b$; however, Bob’s bid is confirmed and Alice’s is not. Now, we can have Alice and Bob swap their bids. The miner creates the same block as before in which the position originally corresponding to Bob now has Alice’s bid of $b^{\prime}$. Since the mechanism is weakly symmetric (Definition 1), Alice’s bid is confirmed. This way, the social welfare increases by $b-b^{\prime}$ in comparison with the honest case, and this violates global SCP. ∎
Lemma 4.2.
For any global SCP mechanism, the amount of burnt coins depends only on the number of confirmed bids.
Proof.
Suppose in two different scenarios, when everyone acts honestly, the blocks made are ${\bf B}$ and ${\bf B}^{\prime}$ respectively, the confirmed bids are ${\mathbf{b}}\subseteq{\bf B}$ and ${\mathbf{b}}^{\prime}\subseteq{\bf B}^{\prime}$ respectively where ${\mathbf{b}}$ and ${\mathbf{b}}^{\prime}$ are of the same length, and the burnt amount in the two scenarios are $q$ and $q^{\prime}$ respectively, where $q<q^{\prime}$. Now, suppose we are actually in the second scenario. A global coalition can adopt the following strategy: create a block identical to ${\bf B}$ in which the confirmed bids correspond to the users with the highest true values and the rest can be fake bids. Observe that the social welfare is the sum of the true values of all confirmed bids (where fake bids have a true value of 00) minus the total coins burnt. Therefore, the above strategy achieves strictly higher social welfare than the honest case. ∎
Theorem 4.3.
No non-trivial deterministic TFM can simultaneously satisfy UIC, MIC, and global SCP when the block size is finite.
Proof.
Since the block size is finite, there must exist a scenario that maximizes the number of confirmed bids. Suppose in this scenario, the block created is ${\bf B}$ and the confirmed bids contained in ${\bf B}$ are ${\mathbf{b}}=(b_{1},\ldots,b_{m})$ sorted from high to low. Since the mechanism is non-trivial, we have $m\geq 1$. Given Lemma 4.2, we can use $q_{i}$ to denote the amount of burnt coins when exactly $i$ bids are confirmed.
First, we show that in any scenario with at least $m$ users whose true values are sufficiently large w.r.t. $q_{0},q_{1},\ldots,q_{m}$, if everyone acts honestly, the mechanism must confirm $m$ bids — by Lemma 4.1, they must also be the $m$ highest bids. Suppose this is not the case, say, $m^{\prime}<m$ bidders are confirmed — by Lemma 4.1, they must be the $m^{\prime}$ highest bids. Now, the global coalition can create the block ${\bf B}$ where the $m$ confirmed bids correspond to the users with the highest true values, and the rest can just be fake bids. In comparison with the honest case, the increase in social welfare is $\sum_{i=m^{\prime}+1}^{m}\left(v_{i}-(q_{i}-q_{i-1})\right)$ which is positive as long as users’ true values $v_{i}$’s are sufficiently large. This violates global SCP.
Imagine a scenario with at least $m+1$ users whose true values are sufficiently high, and $v_{1}\geq\cdots\geq v_{m}>v_{m+1}$. By Lemma 4.1, it must be $v_{1},\dots,v_{m}$ are confirmed. Because the mechanism is UIC, by Myerson’s lemma, all $m$ confirmed bids must pay $v_{m+1}$. However, such a mechanism violates MIC, because the miner can pretend that there is a fake bid $v^{\prime}\in(v_{m+1},v_{m})$. Because the burnt amount is fixed to $q_{m}$, the miner will get more revenue if every confirmed bid pays $v^{\prime}$ instead of $v_{m+1}$. ∎
5 Impossibility of UIC + MIC + Global SCP for Randomized Mechanisms
In this section, we extend the finite-block impossibility of UIC + MIC + global SCP to even randomized mechanisms. Recall that a TFM consists of five rules as defined in Section 2.1, and a randomized TFM may use randomness in any of the five rules. Since the confirmation, the payment, and the miner revenue rules are executed by the blockchain, the strategic players can only bias the randomness in and deviate from the bidding rule and the inclusion rule. Again, due to the revelation principle proven in Section 8, it suffices to consider truthful mechanisms.
5.1 Proof Roadmap
A new challenge in the randomized setting is that whether a bid is confirmed becomes probabilistic, and the arguments in Section 4 no longer hold. Here we give a brief roadmap of our new proof. The key idea of the proof is captured in Lemma 5.4, which says that if there are $i$ equal bids $(\underbrace{b,\ldots,b}_{i})$, followed by other bids $(b_{i+1},\ldots,b_{k+1})$ that are strictly smaller than $b$, then $b_{i+1},\ldots,b_{k+1}$ must be unconfirmed with probability $1$, and moreover, all the $b$-bids must pay $b$ when confirmed, i.e., the total user utility is 00 — throughout, we use $k$ to denote the block size. Since Lemma 5.4 imposes a very strong constraint on the TFM, using it to lead to the final impossibility is not too hard (see Theorem 5.6).
We therefore provide some intuition for the proof of Lemma 5.4. Specifically, we use an inductive proof. The base case is the scenario $(\underbrace{b,\ldots,b}_{k},b_{k+1})$ where $b_{k+1}<b$. Using global SCP and the fact that the block size is only $k$, we show that the bid $b_{k+1}$ must be unconfirmed with probability $1$. Then, we show that the $k$ bids at $b$ must pay $b$ if they are confirmed. To show this, we first argue that in a TFM that is MIC and global SCP, the total user utility is equivalent if $b_{k+1}$ is anything below $b$ (see Lemma 5.3). We then consider a scenario where the $(k+1)$-th bid $b_{k+1}<b$ is arbitrarily close to $b$, and consider lowering one of the first $k$ bids (say, the first bid) to $b_{k+1}-\delta$ — in this case, we argue that the first bid will become unconfirmed using global SCP. Finally, by Myerson’s lemma, we conclude that in the original scenario $(\underbrace{b,\ldots,b}_{k},b_{k+1}<b)$, the first bid must pay $b$ when confirmed.
Once we prove the base case, we proceed by induction for $i=k-1,k-2,\ldots,1$. The induction step is similar in structure to the base case, except that we first need to use the induction hypothesis that in the scenario $(\underbrace{b,\ldots,b}_{i},b_{i+1}<b,\ldots)$, all the $b$ bids pay $b$ when confirmed, to conclude that in the scenario $(\underbrace{b,\ldots,b}_{i-1},b_{i}<b,\ldots)$, $b_{i}$ is unconfirmed with probability $1$ — this follows due to Myerson’s lemma. From this point on, the rest of the proof for the induction step is similar to the base case.
5.2 Formal Proofs
In the rest of this section, we present the formal proofs.
Lemma 5.1.
Suppose the mechanism satisfies global SCP. Then, for any bid vector ${\mathbf{b}}=(b_{1},\dots,b_{t})$, if $b_{i}>b_{j}$, it must be that $x_{i}({\mathbf{b}})\geq x_{j}({\mathbf{b}})$ where $x_{i}({\mathbf{b}})$ and $x_{j}({\mathbf{b}})$ denote user $i$ and $j$’s confirmation probabilities, respectively.
Proof.
For the sake of contradiction, suppose there exists two integers $i,j$ such that $b_{i}>b_{j}$ and $x_{i}({\mathbf{b}})<x_{j}({\mathbf{b}})$. The global coalition can swap user $i$’s and user $j$’s bids. Concretely, user $i$ should bid $b_{j}$ and user $j$ should bid $b_{i}$. Since the bid vector is the same as before (though $b_{i}$ represents user $j$’s bid, and $b_{j}$ represents user $i$’s bid), the expected burning is the same as before. However, since the mechanism is weakly symmetric, the social welfare increases by $\left(b_{i}\cdot x_{j}({\mathbf{b}})+b_{j}\cdot x_{i}({\mathbf{b}})\right)-\left(b_{i}\cdot x_{i}({\mathbf{b}})+b_{j}\cdot x_{j}({\mathbf{b}})\right)>0$, which violates global SCP. ∎
Lemma 5.2.
Suppose the mechanism satisfies global SCP and let the block size be $k$. Consider any bid vector satisfying $b_{1}\geq b_{2}\geq\ldots\geq b_{k}>b_{k+1}$. Then, it must be that $b_{k+1}$ has 0 probability of confirmation.
Proof.
Suppose for the sake of contradiction that $b_{k+1}$ has non-zero probability of confirmation. This means it must be included in the block with some non-zero probability $p$, and conditioned on being included, it has a non-zero probability $p^{\prime}$ of confirmation. A global coalition whose true values are $b_{1},\ldots,b_{k+1}$ can play the following strategy. First, run the honest mechanism in its head including flipping any coins required by the inclusion rule assuming everyone bids honestly, and let ${\bf B}$ be the resulting block. If $b_{k+1}\notin{\bf B}$, just have everyone bid truthfully and build the block ${\bf B}$. Else, if $b_{k+1}\in{\bf B}$, then let $b_{i}>b_{k+1}$ be one bid that is left out of the block — such a bid must exist because the block size is only $k$. The coalition will have the $i$-th user actually bid $b_{k}$ and build the same block ${\bf B}$ where the $b_{k+1}$ actually corresponds to user $i$’s strategic bid. With this strategy, the coalition’s expected gain in utility is $p\cdot p^{\prime}\cdot(b_{i}-b_{k+1})>0$, which violates global SCP. ∎
Lemma 5.3.
Suppose the mechanism satisfies MIC and global SCP. Let ${\bf a}$ be an arbitrary vector of positive length, and let ${\mathbf{b}},{\mathbf{b}}^{\prime}$ arbitrary vectors — ${\mathbf{b}}$ and ${\mathbf{b}}^{\prime}$ may or may not be of the same length, and their lengths are allowed to be 00. Consider two scenarios with bid vectors $({\bf a},{\mathbf{b}})$ and $({\bf a},{\mathbf{b}}^{\prime})$ respectively. Suppose that ${\mathbf{b}}$ and ${\mathbf{b}}^{\prime}$ have 0 confirmation probability in each of the two scenarios, respectively. Then, both scenarios enjoy the same expected miner utility, total social welfare, and total user utility.
Proof.
We first prove that expected miner utility is the same in both scenarios. Suppose this is not true, and without loss of generality, suppose expected miner utility is higher in scenario 1. Then, the miner can ignore the bids ${\mathbf{b}}$, inject the fake bids ${\mathbf{b}}^{\prime}$, pretend that the bid vector is $(\mathbf{a},{\mathbf{b}}^{\prime})$, and run the honest mechanism. Since the confirmation probability of ${\mathbf{b}}^{\prime}$ is 00, the miner need not pay any cost for the fake bids. Therefore, the miner gets higher expected utility by taking the above strategy which violates MIC.
The proof of total social welfare is similar. Suppose, without loss of generality, that the expected total social welfare in scenario 1 is higher. Then, the global coalition can inject fake bids ${\mathbf{b}}^{\prime}$ and pretend that the bid vector is $(\mathbf{a},{\mathbf{b}}^{\prime})$, thus allowing it to increase its expected social welfare. This violates global SCP.
The equivalence in total user utility follows directly from the above, since total user utility is the difference between the social welfare and the miner utility. ∎
Lemma 5.4.
Suppose the mechanism satisfies UIC, MIC, and global SCP, and the block size is $k$. Let ${\mathbf{b}}=(b_{1},\dots,b_{k+1})$ be any vector where $b_{1}=b_{2}=\cdots=b_{i}>b_{i+1}>\cdots>b_{k+1}$. Then, under the bid vector ${\mathbf{b}}$, the following hold: 1) users $i+1,\ldots,k+1$ have 00 confirmation probability; and 2) the total user utility is 00 assuming every one is bidding their true value.
Proof.
We prove both statements by induction. Henceforth, given a bid vector ${\mathbf{b}}=(b_{1},\ldots,b_{k+1})$, we use $x_{i}({\mathbf{b}})$ to denote the confirmation probability of the $i$-th user under ${\mathbf{b}}$.
Base case.
The base case is when $i=k$; that is, there are $k+1$ users with valuations ${\mathbf{b}}=(b_{1},\dots,b_{k+1})$ where $b_{1}=\cdots=b_{k}>b_{k+1}$. By Lemma 5.2, we have $x_{k+1}({\mathbf{b}})=0$. Next, we show that the joint utility of all users must be zero. By Lemma 5.3 and Lemma 5.2, the total user utility under ${\mathbf{b}}=(\underbrace{b,\ldots,b}_{k},b_{k+1})$ is the same as under ${\mathbf{b}}^{\prime}=(\underbrace{b,b,\ldots,b}_{k},b-\delta)$ for any arbitrarily small $\delta>0$ — in both cases, the last user must have 0 confirmation probability. Now, consider the scenario ${\mathbf{b}}^{\prime}=(\underbrace{b,b,\ldots,b}_{k},b-\delta)$. If one of the $b$-bidders denoted $i\in[k]$ lowered its bid to less than $b-\delta$ (and everyone else’s bids stay the same), its confirmation probability would become 00 by Lemma 5.2. Since the mechanism is UIC, by Myerson’s lemma, under $(b,b,\ldots,b,b-\delta)$, user $i$’s expected payment must be at least $(b-\delta)\cdot x_{i}({\mathbf{b}}^{\prime})$. Therefore, user $i$’s expected utility is at most $\delta\cdot x_{i}({\mathbf{b}}^{\prime})$. Therefore, the expected total user utility under ${\mathbf{b}}^{\prime}$ is at most $k\cdot\delta\cdot x_{i}({\mathbf{b}}^{\prime})\leq k\cdot\delta$. This means that under ${\mathbf{b}}=(b_{1},\ldots,b_{k+1})$, the expected total user utility is at most $k\cdot\delta$ for any arbitrarily small $\delta>0$, i.e., the expected total user utility is at most 00. Finally, by the individual rationality of the payment rule, each user’s utility is non-negative when it bids truthfully. Therefore, the expected total user utility is 00.
Inductive step.
Fix any $i\in\{2,\dots,k\}$, suppose the lemma statement holds. We will show that the lemma also holds for any vector ${\mathbf{b}}=(b_{1},\dots,b_{k+1})$ where $b_{1}=b_{2}=\cdots=b_{i-1}>b_{i}>b_{i+1}>\cdots>b_{k+1}$.
We first show that all users $i,\ldots,k+1$ are unconfirmed. Consider the vector $\underbrace{b,b,\ldots,b}_{i},b_{i+1},\ldots,b_{k+1}$. By our induction hypothesis, $i$ users can have non-zero confirmation probability and moreover, the total user utility is 00. Because of individual rationality, each user’s utility must be non-negative, so every user’s utility is also 00. In this case, whenever a bid is confirmed, the payment is equal to the bid (and hence the true value). By UIC and Myerson’s lemma, it means that if the $i$-th user lowers its bid to $b_{i}<b$ (which becomes the scenario ${\mathbf{b}}$ we care about), then user $i$’s confirmation probability becomes 00. By Lemma 5.1, users $i+1,\ldots,k+1$ must have 0 confirmation probability under ${\mathbf{b}}$ too.
Next, we show that the joint utility of all users is zero under ${\mathbf{b}}$ and assuming everyone bids truthfully. By Lemma 5.3, the total user utility under ${\mathbf{b}}$ is the same as under the scenario ${\mathbf{b}}^{\prime}=(\underbrace{b,b,\ldots,b}_{i-1},b-\delta,b-2\delta,\ldots,b-(k+2-i)\delta)$ for an arbitrarily small $\delta>0$ — since in both scenarios, users $i,\ldots,k+1$ have 0 confirmation probability. Under ${\mathbf{b}}^{\prime}$, if any $b$-bidders denoted $j$ lowers its bid to less than $b-(k+2-i)\delta$, its confirmation probability becomes 00 by Lemma 5.2. By UIC and Myerson’s lemma, $j$’s expected payment under ${\mathbf{b}}^{\prime}$ is at least $(b-(k+2-i)\delta)\cdot x_{j}({\mathbf{b}}^{\prime})$. This means that the total user utility under ${\mathbf{b}}^{\prime}$ is at most $(k+2-i)\delta\cdot x_{j}({\mathbf{b}}^{\prime})\cdot(i-1)\leq k^{2}\delta$. Therefore, under ${\mathbf{b}}$, the total user utility is at most $k^{2}\delta$ for any arbitrarily small $\delta>0$. This means that the total user utility under ${\mathbf{b}}$ is at most 00. Finally, by the individual rationality of the payment rule, each user’s utility is non-negative when it bids truthfully. Therefore, the expected total user utility is 00.
Lemma 5.5.
Suppose the mechanism satisfies UIC, MIC, and global SCP, and the block size is $k$. Let $a$ be any positive real number. Consider a scenario with only one bid $a$. Then, the only user’s utility is zero assuming it bids its true value.
Proof.
Fix $a_{1}>0$. Consider any vector $\mathbf{a}=(a_{1},a_{2},\dots,a_{k+1})$ such that $a_{1}>a_{2}>\cdots>a_{k+1}$. By Lemma 5.4, under $\mathbf{a}$, all users except the first user must have 0 confirmation probability, and moreover, the first user has utility 00. By Lemma 5.3, the total expected user utility under $\mathbf{a}$ and under a single bid $a_{1}$ are the same. Therefore, under a single bid $a_{1}$, the user’s utility is also 00. ∎
Theorem 5.6.
No non-trivial, possibly randomized TFM can simultaneously satisfy UIC, MIC, and global SCP when the block size is finite.
Proof.
We will show that under any sufficiently large $a$, the confirmation probability under a single bid $a$ is non-zero. If we can show this, then we can show a contradiction to UIC. Specifically, consider $b>a$ and both sufficiently large. By Lemma 5.5, if there is only one user with true value $b$, its utility is zero when it bids truthfully. However, the user can underbid $a$. Since the confirmation probability is non-zero and the payment is at most $a$, the user enjoys positive utility, which violates UIC.
We now show that the confirmation probability under a single bid $a$ for a sufficiently large $a$ is non-zero assuming the miner honestly follows the inclusion rule. Suppose this is not true for the sake of a contradiction; that is, when the bid vector is just $a$ and the miner is honest, no one has positive confirmation probability, so the social welfare is zero. Since the mechanism is non-trivial, there must exist a scenario ${\mathbf{b}}=(b_{1},\dots,b_{t})$ such that some user $i$ has positive confirmation probability. Let $a>\sum_{j=1}^{t}b_{j}$. Then, imagine that there is only one user with true value $a$. The global coalition can replace the user’s primary bid with $b_{i}$, inject the fake bids ${\mathbf{b}}_{-i}$, and pretend that the bid vector is ${\mathbf{b}}$. This way, the user would get positive confirmation probability and since its payment is at most $\sum_{j=1}^{t}b_{j}$, the user’s expected utility (and thus the social welfare) is positive. This violates global SCP.
6 Feasibility and Impossibility of UIC + MIC + OCA-Proof
We can generalize the proof in Section 5, and rule out UIC, MIC, and OCA-proof (rather than global SCP) for truthful mechanisms. Recall that for a truthful mechanism, the difference between OCA-proof and global SCP is that global SCP insists that the optimal strategy of the global coalition is the truthful strategy, whereas OCA-proofness allows it to be some other strategy in which each user acts independently and bids the outcome of some function $\sigma(\cdot)$.
Interestingly, if we allow the bidding rule to be not truth-telling, i.e. considering non-truthful mechanisms, we can have a mechanism that satisfies UIC, MIC, and OCA-proof. We present the feasibility for non-truthful mechanisms in Section 6.1, and we prove the impossibility of UIC + MIC + OCA-proof for truthful mechanisms in Section 6.2. Notice that because of the feasibility in Section 6.1, we must require the bidding rule to be truth-telling to reach an impossibility in Section 6.2.
6.1 A Non-Truthful Mechanism with UIC + MIC + OCA-Proof
The rationale of the design is to signal to the mechanism when everyone is adopting the globally optimal strategy $\sigma$ (as opposed to the bidding rule used to establish UIC). When the mechanism detects that everyone is behaving according to $\sigma$, it adopts a different behavior to optimize social welfare. We use the range $[0,1)$ to encode the actual bid, and use the range $[1,\infty)$ for signalling. While the resulting mechanism is somewhat contrived and not necessarily meaningful from a practical point of view, it clarifies which notions of collusion-resilience most accurately capture the intended modeling goals and illustrates some technical challenges involved in the proof in Section 6.2. Consider the following TFM:
  • Globally optimal strategy $\sigma(v)$: Given a true value $v$, output a bid $v+1$.
  • Bidding rule: Given a true value $v$, output a bid $1/(v+2)$.
  • Inclusion rule: Let $S$ be the set of all pending bids that are in $[0,1)$. If $|S|>k$, then randomly select $k$ bids from $S$ to include. If $1\leq|S|\leq k$, then include all bids in $S$. If $|S|=0$, choose the top up to $k$ bids to include.
  • Confirmation, payment, and miner revenue rules: All included bids are confirmed. Each confirmed bid pays nothing, and the miner gets nothing.
Obviously, this mechanism is non-trivial.
Claim 6.1.
The above mechanism satisfies UIC, MIC, and OCA-proofness.
Proof.
For UIC, notice that if a user follows the bidding rule, its bid is always in $[0,1)$. If there is no bid in $[0,1)$ before a user submits its bid, then bidding $1/(v+2)$ always guarantees user’s bid to be included and confirmed, where $v$ denote the true value. If there is already some bids in $[0,1)$ before a user submits its bid, then bidding $1/(v+2)$ is a dominant strategy since it guarantees the user’s bid is added to $S$, the set of all bids in $[0,1)$, which is the best a user can do. Next, MIC holds since the miner revenue is always zero. Finally, if all users follow the globally optimal strategy $\sigma$, everyone’s bid is at least $1$. The honest inclusion rule will include the top up to $k$ bids, which maximizes the social welfare. Thus, OCA-proofness holds. ∎
Remark 1.
We can try to apply revelation principle, and bake the bidding rule into the mechanism so that the resulting mechanism is truthful. For example, whenever seeing a bid $b$, the miner and the mechanism view it as $1/(b+2)$. The modified mechanism, however, does not satisfy OCA-proofness anymore when the number of users is larger than the block size, since the miner should choose $k$ users with the highest true values instead of the random selection as indicated by the inclusion rule. This is not a coincidence: in the next section, we show that it is impossible to have a non-trivial truthful mechanism satisfying UIC, MIC, and OCA-proofness.
6.2 Impossibility of UIC + MIC + OCA-Proof for Truthful Mechanisms
The structure of the proof in this section is similar to the proof in Section 5. In fact, Lemmas 6.2, 6.3, 6.6, 6.7, 6.8 and 6.9 are the analogs of Lemmas 5.1, 5.2, 5.3, 5.4, 5.5 and 5.6, respectively, except that we need to work on the images of $\sigma$ in order to apply OCA-proofness. Before diving into the proof, we make a few remarks.
  • A key step in Section 5 is to prove a user with true value $v$ must have zero utility, and we prove it by making an unconfirmed bid arbitrarily close to $v$ (see Lemma 5.4). To extend the similar idea to OCA-proofness, we require $\sigma$ to be strictly increasing and continuous. In fact, we show that being strictly increasing is a consequence of UIC and OCA-proofness (Lemma 6.4), and being strictly increasing implies that there exists a point at which $\sigma$ is continuous (Corollary 6.5).
  • As we have seen a feasibility for non-truthful mechanisms in Section 6.1, we must require the bidding rule to be truth-telling to reach an impossibility. Indeed, the proofs of Lemmas 6.4, 6.7, 6.8 and 6.9 rely on Myerson’s lemma, and Myerson’s lemma requires the mechanism to be truthful. Notice that the mechanism in Section 6.1 does not satisfy the requirements of Myerson’s lemma. If we apply the revelation principle to make the mechanism become truthful so that Myerson’s lemma could apply, it breaks OCA-proofness (see Remark 1).
  • Notice that OCA-proofness requires $\sigma$ to output a single real-valued bid. In fact, if we allow $\sigma$ to output multiple bids, we can have a (somewhat contrived) mechanism that satisfies UIC, MIC, and OCA-proof (see Section 7.2). In particular, Lemma 6.4 requires $\sigma$ to only output a single real number.
  • The original definition of OCA-proofness in [Rou21] also requires $\sigma$ to be individually rational, but this will not be needed for the impossibility proof.
Given any vector ${\mathbf{b}}=(b_{1},\dots,b_{t})$ of the length $t$, let $\sigma({\mathbf{b}})$ denote the element-wise application of $\sigma$, i.e. $\sigma({\mathbf{b}})=\left(\sigma(b_{1}),\dots,\sigma(b_{t})\right)$.
Lemma 6.2.
Suppose the mechanism satisfies OCA-proofness with the globally optimal strategy $\sigma$. Then, for any bid vector $\sigma({\mathbf{b}})=(\sigma(b_{1}),\dots,\sigma(b_{t}))$, if $b_{i}>b_{j}$, it must be that the confirmation probability of the bid $\sigma(b_{i})$ is at least as high as that of $\sigma(b_{j})$.
Proof.
Imagine that $t$ players each have true values $b_{1},\ldots,b_{t}$, and they all bid according to the globally optimal strategy $\sigma$. Let $x_{i},x_{j}$ denote the confirmation probabilities of $\sigma(b_{i})$ and $\sigma(b_{j})$ respectively. Suppose for the sake of contradiction that $b_{i}>b_{j}$, and $x_{j}>x_{i}$. The global coalition can engage in the following OCA: have users $i$ and $j$ swap their bids, that is, user $i$ who has true value $b_{i}$ bids $\sigma(b_{j})$, and user $j$ who has true value $b_{j}$ bids $\sigma(b_{i})$. The set of bids is the same as before, therefore, the expected burning amount does not change. The expected social welfare thus increases by $b_{i}x_{j}+b_{j}x_{i}-(b_{i}x_{i}+b_{j}x_{j})=(b_{i}-b_{j})(x_{j}-x_{i})>0$. This violates OCA-proofness and the fact that $\sigma$ is the social-welfare-maximizing-strategy of the global coalition. ∎
Lemma 6.3.
Suppose the mechanism satisfies OCA-proofness with the globally optimal strategy $\sigma$, and let the block size be $k$. Given $b_{1}\geq b_{2}\geq\ldots\geq b_{k}>b_{k+1}$, then, under the bid vector $\sigma(b_{1}),\sigma(b_{2}),\ldots,\sigma(b_{k}),\sigma(b_{k+1})$, it must be that $\sigma(b_{k+1})$ has 0 probability of confirmation.
Proof.
Imagine that $k+1$ players each have true values $b_{1},\ldots,b_{k+1}$, and they all bid according to the globally optimal strategy $\sigma$. It suffices to show the following: suppose $\sigma(b_{k+1})$ has a non-zero probability $p$ of being included in the block, then conditioned on being included, its probability of confirmation must be 00. Suppose this is not true, and conditioned on being included, $\sigma(b_{k+1})$ has probability $p^{\prime}>0$ of being confirmed. Then, the global coalition can adopt the following OCA: run the honest inclusion rule in its head, including flipping any coins required by the inclusion rule, and let ${\bf B}$ denote the resulting block. If $\sigma(b_{k+1})$ is not included in ${\bf B}$, simply output the block ${\bf B}$. Else if $\sigma(b_{k+1})$ is included in ${\bf B}$, then let $\sigma(b_{i})$ be some bid that is left out — such a bid must exist because the block size is only $k$. Now, have the $i$-th user bid $\sigma(b_{k+1})$ and make an actual block ${\bf B}$ where the bid $\sigma(b_{k+1})$ corresponds to user $i$’s bid. The expected increase in social welfare relative to everyone adopting $\sigma$ is $p\cdot p^{\prime}\cdot(b_{i}-b_{k+1})>0$, which violates OCA-proofness and the fact that $\sigma$ is the social-welfare-maximizing strategy. ∎
Lemma 6.4.
Suppose a truthful mechanism satisfies UIC and OCA-proofness with the globally optimal strategy $\sigma$, and let the block size be $k$. Then, there exists a constant $c^{*}$ such that for any real numbers $z,z^{\prime}$ satisfying $z>z^{\prime}>c^{*}$, it must be $\sigma(z)>\sigma(z^{\prime})$. Conceptually, $\sigma$ must be strictly increasing for large inputs.
Proof.
Since the mechanism is non-trivial, there exists a bid vector $(b_{1},\dots,b_{t})$ such that the total confirmation probability is positive. Let $b_{i}$ be the bid with the highest confirmation probability among $b_{1},\dots,b_{t}$. Let $x$ be the confirmation probability of $b_{i}$ and let $c^{*}$ be any real number larger than $\sum_{j=1}^{t}b_{j}/x$. We will show that for any $z>z^{\prime}>c^{*}$, $\sigma(z)>\sigma(z^{\prime})$.
Consider a scenario where there are $k+1$ users all with the true value $z$. Suppose they form a global coalition and follow the bidding strategy $\sigma$ so that the bid vector is $\mathbf{s}=(\underbrace{\sigma(z),\dots,\sigma(z)}_{k+1})$. There are two possible cases.
First, there exists a user $j$ with positive confirmation probability assuming the miner honestly follows the inclusion rule and select from the bid vector $\mathbf{s}$. Because the mechanism satisfies UIC, by Myerson’s lemma, if user $j$ increases its bid, the confirmation probability must be non-decreasing. Thus, if $\sigma(z)\leq\sigma(z^{\prime})$, and user $j$ increases its bid from $\sigma(z)$ to $\sigma(z^{\prime})$, its confirmation probability is still positive. However, if user $j$ bids $\sigma(z^{\prime})$, the resulting bid vector will become $(\underbrace{\sigma(z),\dots,\sigma(z)}_{k},\sigma(z^{\prime}))$. By Lemma 6.3, the bid $\sigma(z^{\prime})$ has zero confirmation probability, which leads to a contradiction. Consequently, it must be $\sigma(z)>\sigma(z^{\prime})$.
Second, if everyone’s confirmation probability is zero assuming the miner honestly follows the inclusion rule and select from the bid vector $\mathbf{s}$, following $\sigma$ will lead to zero social welfare. We argue that this case is impossible by showing an alternative strategy of the global coalition that achieves positive social welfare, contradicting the fact that $\sigma$ is social-welfare-maximizing. Specifically, the alternative strategy works as follows: the global coalition pretends that the bid vector is $(b_{1},\dots,b_{t})$ where the bid $b_{i}$ corresponds to the primary bid of any user, say, the first user. Recall that the first user’s true value is $z$. Thus, social welfare is at least $z\cdot x-p$, where $p$ is the expected payment of all users. Since $b_{i}$ has the highest confirmation probability among $b_{1},\dots,b_{t}$, $p$ is at most $x\cdot\sum_{j=1}^{t}b_{j}$. Because $z>c^{*}>\sum_{j=1}^{t}b_{j}/x$, the social welfare is positive.
Corollary 6.5.
Suppose a truthful mechanism satisfies UIC and OCA-proofness with the globally optimal strategy $\sigma$, and let the block size be $k$. Then, there exists a constant $c^{*}$ such that for any real number $a>c^{*}$, there exists a real number $a^{*}\geq a$ at which $\sigma$ is continuous.
Proof.
A well-known fact is that any function that is monotone must have only countably many discontinuities within any interval $[x_{1},x_{2}]$ where $x_{2}>x_{1}$; further, all such discontinuities must be jump discontinuities. By Lemma 6.4, there exists a constant $c^{*}$ such that $\sigma$ is monotone for all inputs larger than $c^{*}$. Thus, for any $a>c^{*}$, within the interval $[a,a+1]$, $\sigma$ is monotone and there must exist a point $a^{*}$ such that $\sigma$ is continuous at $a^{*}$. ∎
Lemma 6.6.
Suppose the mechanism satisfies MIC and OCA-proofness with the globally optimal strategy $\sigma$. Let ${\bf a},{\mathbf{b}},{\mathbf{b}}^{\prime}$ be arbitrary vectors. Consider two scenarios $(\sigma({\bf a}),\sigma({\mathbf{b}}))$ and $(\sigma({\bf a}),\sigma({\mathbf{b}}^{\prime}))$. Suppose that $\sigma({\mathbf{b}})$ and $\sigma({\mathbf{b}}^{\prime})$ have 0 confirmation probability in each of the two scenarios, respectively. Then, both scenarios enjoy the same expected miner utility, social welfare, and total user utility.
Proof.
The equivalence in expected miner utility follows directly from MIC and the fact that the miner can change from ${\mathbf{b}}$ to ${\mathbf{b}}^{\prime}$ for free, or vice versa, since these bids are unconfirmed.
The equivalence in social welfare follows from OCA-proofness, and the fact that $\sigma$ maximizes social welfare. Suppose, for example, the social welfare under $(\sigma(\mathbf{a}),\sigma({\mathbf{b}}))$ is strictly greater than $(\sigma(\mathbf{a}),\sigma({\mathbf{b}}^{\prime}))$. Then, under the scenario where the users’ true values are $(\mathbf{a},{\mathbf{b}}^{\prime})$, everyone following $\sigma$ will not maximize social welfare, since there is an OCA that have the users bid $(\sigma(\mathbf{a}),\sigma({\mathbf{b}}))$ instead which strictly increases the social welfare.
The equivalence in total user utility follows directly from the above, since total user utility is the difference between the social welfare and the miner utility. ∎
Lemma 6.7.
Suppose a truthful mechanism satisfies UIC, MIC and OCA-proofness with the globally optimal strategy $\sigma$, and the block size is $k$. Then, there exists a constant $c^{*}$ such that for any $i\in[k]$, the following holds.
  • Consider a scenario with the arbitrary bid vector $\mathbf{s}=\underbrace{\sigma(b),\ldots,\sigma(b)}_{i},\sigma(b_{i+1}),\ldots,\sigma(b_{k+1})$, where $b>b_{i+1}>b_{i+2}>\ldots>b_{k+1}>c^{*}$. Then, if $\sigma$ is continuous at $b$, the bids $\sigma(b_{i+1}),\ldots,\sigma(b_{k+1})$ have 0 confirmation probability.
  • Moreover, if there are $k+1$ users with the true value $\mathbf{s}$ and $\sigma$ is continuous at $b$, the expected total user utility is 00 when all users bid truthfully.
Proof.
By Lemma 6.4, there exists a constant $c^{*}$ such that for any real numbers $z,z^{\prime}$ satisfying $z>z^{\prime}>c^{*}$, it must be $\sigma(z)>\sigma(z^{\prime})$. Throughout this proof, we assume $b_{k+1}>c^{*}$ so that $\sigma$ is strictly increasing for any input larger than $b_{k+1}$. By Corollary 6.5, there exists a real number $b>b_{k+1}$ such that $\sigma$ is continuous at $b$, and we choose $b$ such that this condition holds. We will prove the lemma by induction.
Base case.
The base case is when $i=k$; that is, there are $k+1$ users with the true values $\mathbf{s}=(\underbrace{\sigma(b),\ldots,\sigma(b)}_{k},\sigma(b_{k+1}))$. Notice the true value is $\sigma(b)$ instead of $b$ for the first $k$ users. By Lemma 6.3, $\sigma(b_{k+1})$ has 0 confirmation probability. By Lemma 6.6, the total user utility under $\mathbf{s}$ is the same as under $\mathbf{s}^{\prime}=(\underbrace{\sigma(b),\ldots,\sigma(b)}_{k},\sigma(b-\delta))$ where $\delta>0$ can be arbitrarily small — since in both scenarios, the last bid is unconfirmed. Now, suppose the bid vector is $\mathbf{s}^{\prime}$. Fix any user with the true value $\sigma(b)$, say the first user. If it lowered its bid to anything less than $\sigma(b-\delta)$, its confirmation probability must be 00 by Lemma 6.3. Since the mechanism is UIC, by Myerson’s lemma, the payment of the first user is at least $\sigma(b-\delta)\cdot x_{1}(\mathbf{s}^{\prime})$. In this case, the first user’s utility is at most $(\sigma(b)-\sigma(b-\delta))\cdot x_{1}(\mathbf{s}^{\prime})\leq\sigma(b)-\sigma(b-\delta)$. Therefore, the total user utility under $\mathbf{s}^{\prime}$ (and hence under $\mathbf{s}$) is at most $k\cdot(\sigma(b)-\sigma(b-\delta))$. Since $\sigma$ is continuous at $b$, and the above holds for any $\delta>0$, the total user utility under $\mathbf{s}$ is at most 00. By the individual rationality of the payment rule, each user’s utility is non-negative when it bids truthfully. Therefore, the expected total user utility is 00.
Inductive step.
Fix any $i\in\{2,\dots,k\}$, and suppose the lemma holds. We want to show that the lemma also holds for any $\mathbf{s}_{i-1}=(\underbrace{\sigma(b),\ldots,\sigma(b)}_{i-1},\sigma(b_{i}),\sigma(b_{i+1}),\ldots,\sigma(b_{k+1}))$. Start from the scenario $\mathbf{s}_{i}=(\underbrace{\sigma(b),\ldots,\sigma(b)}_{i},\sigma(b_{i+1}),\ldots,\sigma(b_{k+1}))$, and imagine that the $i$-th bid lowers from $\sigma(b)$ to $\sigma(b_{i})$. Since the $i$-th user has 0 utility under $\mathbf{s}_{i}$, by UIC and Myerson’s lemma, its confirmation probability must become 0 when it lowers its bid to $\sigma(b_{i})<\sigma(b)$. By Lemma 6.2, the bids $\sigma(b_{i+1}),\ldots,\sigma(b_{k+1})$ must have 0 confirmation probability too.
We next show that under $\mathbf{s}_{i-1}$, total user utility is 00. Consider another bid vector
$\mathbf{s}^{\prime}:=(\underbrace{\sigma(b),\ldots,\sigma(b)}_{i-1},\sigma(b-\delta),\ldots,\sigma(b-(k+2-i)\delta))$ for some $\delta>0$. By the argument above, the bids numbered $i+1,\ldots,k+1$ are not confirmed. By Lemma 6.6, the total user utility under $\mathbf{s}_{i-1}$ is the same as under the scenario $\mathbf{s}^{\prime}$. In scenario $\mathbf{s}^{\prime}$, if any of the $\sigma(b)$ bidders, say the first bidder, lowers its bid from $\sigma(b)$ to anything less than $\sigma(b-(k+2-i)\delta)$, its confirmation probability becomes 00 by Lemma 6.3. By UIC and Myerson, the first user’s payment under $\mathbf{s}^{\prime}$ is at least $\sigma(b-(k+2-i)\delta)\cdot x_{1}(\mathbf{s}^{\prime})$, and its utility is at most $(\sigma(b)-\sigma(b-(k+2-i)\delta))\cdot x_{1}(\mathbf{s}^{\prime})\leq\sigma(b)-\sigma(b-(k+2-i))\delta$. Therefore, the total user utility under $\mathbf{s}^{\prime}$ (and also under $\mathbf{s}_{i-1}$) is upper bounded by $k\cdot(\sigma(b)-\sigma(b-(k+2-i)\delta))$. Since $\sigma$ is continuous at $b$ and the above holds for any arbitrarily small $\delta>0$, it means that the total user utility under $\mathbf{s}_{i-1}$ is at most 00. By the individual rationality of the payment rule, each user’s utility is non-negative when it bids truthfully. Therefore, the expected total user utility is 00. ∎
Lemma 6.8.
Suppose a truthful mechanism satisfies UIC, MIC and OCA-proofness with the globally optimal strategy $\sigma$, and the block size is $k$. Then, there exists a constant $c^{*}$ such that for any $a^{*}>c^{*}$, if there is only a single user with the true value $\sigma(a^{*})$ and $\sigma$ is continuous at $a^{*}$, its utility is zero when it bids truthfully.
Proof.
By Lemma 6.7, there exists a constant $c^{*}$ such that under the scenario $\mathbf{a}:=\left(\sigma(a^{*}),\sigma(a_{2}),\ldots,\sigma(a_{k+1})\right)$ where $a^{*}>a_{2}>\ldots>a_{k+1}>c^{*}$, the utility of the first user is 00 assuming it is bidding truthfully. Moreover, all users except the first user must have 0 confirmation probability. By Lemma 6.6, the total expected user utility under $\mathbf{a}$ and under a single bid $\sigma(a^{*})$ are the same. Therefore, under a single bid $\sigma(a^{*})$, the user’s utility is also 0. ∎
Theorem 6.9.
No non-trivial, possibly randomized truthful TFM can simultaneously satisfy UIC, MIC, and OCA-proofness when the block size is finite.
Proof.
We will show that under any sufficiently large $a$ at which $\sigma$ is continuous, the confirmation probability under a single bid $\sigma(a)$ is non-zero. If we can show this, then we can show a contradiction to UIC and the Myerson’s lemma. Specifically, consider $b>a$ which both are sufficiently large and at which $\sigma$ is continuous (Corollary 6.5 guarantees such $a,b$ exist). By Lemma 6.8, under both scenarios, the user utility is 00 (assuming the bid, $\sigma(a)$ or $\sigma(b)$, is the same as the user’s true value). By Lemma 6.4, for sufficiently large $a$, we have $\sigma(b)>\sigma(a)$. However, by Myerson’s lemma, if the confirmation probabilities under a single $\sigma(a)$ and under a single $\sigma(b)$ are both non-zero, it cannot be that in both cases the user’s utility is 00. Otherwise, the user with the true value $\sigma(b)$ can underbid $\sigma(a)$. Since the confirmation probability is non-zero and the payment is at most $\sigma(a)$, the user enjoys positive utility, which violates UIC.
We now show that the confirmation probability under a single bid $\sigma(a)$ for a sufficiently large $a$ is non-zero assuming the miner honestly follows the inclusion rule. Suppose this is not true for the sake of a contradiction; that is, when the bid vector is just $\sigma(a)$ and the miner is honest, no one has positive confirmation probability, so the social welfare is zero. Since the mechanism is non-trivial, there exists a scenario ${\mathbf{b}}=(b_{1},\ldots,b_{t})$ such that the total confirmation probability is non-zero, and let $x>0$ be the confirmation probability of some bid $b_{i}$ in ${\mathbf{b}}$. Consider $a>\sum_{j\in[n]}b_{j}/x$, and suppose there is only one user with the true value $a$. We will show that the strategy $\sigma(a)$ does not maximize social welfare which contradicts to the fact that $\sigma$ is social-welfare-maximizing-strategy. Specifically, the user can post the bids $(b_{1},\ldots,b_{t})$ instead where position $i$ corresponds to the user’s primary bid and all others are fake bids. Note that under ${\mathbf{b}}$, the total user payment is at most $\sum_{j\in[n]}b_{j}$. Since the user’s true value is greater than $\sum_{j\in[n]}b_{j}/x$ and its confirmation probability is $x$, the total user utility (and thus the social welfare) is positive under this new bidding strategy.
Remark 2.
Notice that Chung and Shi [CS23] showed that no TFM can simultaneously satisfy UIC and $1$-SCP when the block size is finite. Their impossibility does not rely on MIC, while MIC is necessary for proving Theorem 6.9. Specifically, consider the second-price auction without burning defined as follows. Given the block size $k$, the miner includes the top up to $k$ bids in the block. If the number of included bids is less than $k$, then all included bids are confirmed, and pay nothing. If the number of included bids is $k$, then the top $k-1$ bids are confirmed, while the $k$-th bid is unconfirmed (break ties arbitrarily). All confirmed bids pay the $k$-th bid, and all the payments go to the miner. This mechanism satisfies UIC since the confirmation and the payment satisfy the requirements of Myerson’s lemma. The mechanism also satisfies OCA-proofness when $\sigma$ is just bidding truthfully since the confirmed bids are always the ones with the highest true values, which maximizes the social welfare. However, it does not satisfy MIC, since the miner is incentivized to inject a fake bid to increase the $k$-th bid.
7 How to Circumvent the Impossibilities
In this section, we discuss possible ways to circumvent the impossibilities. Recall that in the definition of OCA-proofness, we require all users act independently in the globally optimal strategy $\sigma$, and $\sigma$ should only output a single bid. If we remove either of the two requirements, we can have mechanisms that satisfies UIC, MIC and OCA-proofness, where we present in Sections 7.1 and 7.2. These mechanisms are somewhat contrived and not obviously relevant for real-world blockchain protocols; the primary point of these mechanisms is to “stress-test” the definitions, and these examples highlight the subtlety of the modeling and help us to have a better understanding of the trade-offs between various notions of collusion-resilience. In Section 7.3, we consider an even weaker definition, called inclusion-rule-respecting, which might be the weakest but still meaningful definition that captures the notion of “no way to steal from the protocol.” Finally, in Section 7.4, we discuss the use of cryptography in the TFM literature, and point out a future direction by using cryptography and Bayesian incentive compatibility.
7.1 Allowing the Globally Optimal Strategy to Coordinate
OCA-proofness requires that all users act independently in the globally optimal strategy $\sigma$. However, if we remove this restriction, then the following auction would simultaneously satisfy UIC, MIC, and OCA-proof (with the aforementioned relaxation).
Posted price auction with random selection and total burning.
The mechanism is parametrized with some reserve price $r>0$ and block size $k$. Any bid that is at least $r$ is considered eligible. Randomly select up to $k$ bids to include in the block. All included bids are confirmed and they each pay $r$. All payment is burnt and the miner receives nothing. It is not hard to see that the above mechanism satisfies UIC and MIC. It also satisfies OCA-proofness with the aforementioned relaxation since the global coalition can just have the users with the top $k$ values bid more than $r$, while everyone else bids 00.
Note that allowing all users to coordinate bids does not trivialize the definition, as this relaxed OCA-proofness definition still requires that the miner follow the TFM’s inclusion rule honestly. For example, in the auction above, the miner has to select $k$ bids uniformly at random among all eligible bids.
7.2 Allowing the Globally Optimal Strategy to Output Multiple Bids
We next show that when we allow the globally optimal strategy $\boldsymbol{\sigma}$ to output multiple bids, then we can actually construct a truthful mechanism that satisfies UIC, MIC, and OCA-proofness. Similar to Section 6.1, we try to signal to the mechanism when everyone is adopting the globally optimal strategy $\sigma$. In this mechanism, since $\sigma$ can output multiple bids, users can signal through fake bids. We assume the miner and the mechanism can distinguish whether a set of bids come from the same user. In practice, this can be implemented by checking whether these bids are signed by the same public key. Specifically, consider the following mechanism parametrized by a reserve price $r>1$.
  • Globally optimal strategy $\boldsymbol{\sigma}(v)$: If the user’s true value $v\geq r$, $\boldsymbol{\sigma}(v)$ outputs $(r,r/v)$ where $r/v\in(0,1]$ is a uniquely decodable encoding of the user’s true value $v$. Otherwise, if $v<r$, $\boldsymbol{\sigma}(v)$ simply outputs $\emptyset$, i.e., the user drops its bid.
  • Inclusion rule.
    • Case 1: Check if it is the case that every user $i\in[n]$ submitted a message of the form $(r,r/v_{i})$ where $r/v_{i}\in(0,1]$ for all $i\in[n]$. If so, decode the true value vector $(v_{1},\ldots,v_{n})$. Choose the $\min(n,k)$ bids with the highest true values, and include their corresponding $r$ bids in the block.
    • Case 2: Else, if the number of bids is even, henceforth denoted $2n$, check whether there are exactly $n$ bids at $r$ and $n$ number of bids that lie in the range $(0,1]$. If so, interpret the bid vector as $(r,r/v_{1}),\ldots,(r,r/v_{n})$, as if each pair $(r,r/v_{i})$ comes from the $i$-th user for $i\in[n]$. Now, apply the algorithm of Case 1.
    • Case 3: Else, if the number of bids is odd, henceforth denoted $12n+$ 1, check if there are $n$ bids at $r$ and $n$ bids in the range of $(0,1]$, and one last bid denoted $r^{\prime}$. If $r^{\prime}>r$, then choose $r^{\prime}$ to include in the block, and among the remaining $n$ bids at $r$, randomly choose $\min(k-1,n)$ of them to include in the block; else if $r^{\prime}\leq r$, randomly choose $\min(k,n)$ bids among all the $r$-bids to include in the block.
    • Case 4: In all other cases, randomly choose $\min(k,n)$ bids among all bids that are strictly larger than $r$ to include in the block.
  • Confirmation, payment, and miner revenue rules: All included bids that are at least $r$ are confirmed. Each confirmed bid pays $r$, and the miner gets nothing.
Claim 7.1.
The above mechanism satisfies UIC, MIC, and a relaxed notion of OCA-proofness in which $\sigma$ is allowed to output multiple bids.
Proof.
MIC is obvious since the miner always gets 0 revenue. For OCA-proofness, observe that if the global coalition all take the strategy $\boldsymbol{\sigma}$, then case 1 of the inclusion rule will be triggered, and thus the social welfare is maximized.
Proving UIC is the most complicated part. Fix any user $i$. Notice that each confirmed bid pays $r$, so if user $i$’s true value is at most $r$, it cannot get positive utility regardless its bid. Thus, bidding truthfully is a dominant strategy. Henceforth, we assume user $i$’s true value is strictly larger than $r$. There are two possibilities.
  1. 1.
    First, before user $i$ submits its bid, the bid vector contains exactly $n$ bids at $r$ and $n$ bids lie in the range $(0,1]$ for some $n$. Then, bidding truthfully will lead to Case 3, where user $i$’s bid is guaranteed to be included and confirmed, and thus maximizes the utility.
  2. 2.
    Second, before user $i$ submits its bid, the bid vector does not contain exactly $n$ bids at $r$ and $n$ bids lie in the range $(0,1]$ for some $n$. Suppose there is no bid strictly larger than $r$ in the bid vector before user $i$ submits its bid. Then, if user $i$ bids truthfully, its bid will be the only bid larger than $r$, and the inclusion rule goes to Case 4. Since $k\geq 1$, user $i$’s bid is guaranteed to be included and confirmed, and thus maximizes the utility. On the other hand, suppose there are $t\geq 1$ bids strictly larger than $r$ in the bid vector before user $i$ submits its bid. Then, the inclusion can never go to Case 1 or Case 2. If user $i$ bids truthfully, the inclusion rule must go to Case 4, and user $i$’s bid is included with probability $\min(k,t+1)/(t+1)$. However, if the strategic bidding leads to Case 3, it must be $t=1$. Then, user $i$’s inclusion probability has dropped from $1/2$ to 00 if $k=1$, or from $1$ to at most 1 if $k\geq 2$. On the other hand, if the strategic bidding leads to Case 4, any untruthful bidding $>r$ leads to the same inclusion probability and the same payment, so the utility does not change, while any untruthful bidding $\leq r$ leads to the zero inclusion probability which leads to zero utility. Finally, notice that injecting fake bids never helps in Case 4. Thus, in all cases, the utility is better than bidding truthfully.
Recall that weakly symmetric mechanisms do not make use of the bidders’ identities or other auxiliary information except for tie-breaking among equal bids. Because the mechanism above relies on the fact that the inclusion rule can distinguish whether a set of bids come from the same user, it may not be obvious that the mechanism satisfies weak symmetry. The next claim proves this formally.
Claim 7.2.
The above mechanism satisfies weak symmetry.
Proof.
To show weak symmetry, we give an equivalent description as follows. Given a bid vector ${\mathbf{b}}$, the sorting algorithm checks whether ${\mathbf{b}}$ satisfies the condition of Case 1. If so, the sorting algorithm places the multiple $r$ bids corresponding to the true value vector $(v_{1},\dots,v_{n})$ (defined in Case 1 above) from high to low, and sorts the rest in the descending order. Otherwise, if ${\mathbf{b}}$ does not satisfy the condition of Case 1, the sorting algorithm sorts ${\mathbf{b}}$ in the descending order and breaks tie arbitrarily.
In the original mechanism, the only case that depends on the extra information (identity, in this case) is Case 1. By applying the sorting algorithm described above, the inclusion rule can implement Case 1 by only depending on the amount of the bids and their relative position in the sorted bid vector.
7.3 Inclusion-Rule-Respecting
It is also natural to consider a further relaxation of OCA-proofness henceforth called “inclusion-rule-respecting”. In comparison with OCA-proofness, inclusion-rule-respecting only requires that the globally optimal strategy $\sigma$ not involve altering the inclusion rule, and all other constraints on $\sigma$ are removed. Inclusion-rule-respecting is strictly weaker than the notions in Section 7.1 and Section 7.2. Therefore, the mechanisms mentioned in Section 7.1 and Section 7.2 simultaneously satisfy UIC, MIC, and inclusion-rule-respecting too.
The “inclusion-rule-respecting” notion has an intuitive interpretation: it discourages miners from altering the intended protocol implementation; in other words, all profiting strategies can be implemented as bidding strategies above the protocol layer. In this sense, “inclusion-rule-respecting” also captures the intuition of “no way to steal from the protocol”. One can also view it as follows: global SCP captures “not stealing from the protocol” where the protocol is the union of the miner’s inclusion rule as well as users’ honest bidding strategies; and inclusion-rule-respecting captures the same notion but where the protocol is only the underlying blockchain protocol. OCA-proofness is somewhere in between.
7.4 Discussions and Open Questions Regarding the Use of Cryptography
Another interesting question is whether we can overcome the impossibilities using cryptography. Shi et al. [SCW23] showed that if the rules of the TFM are enforced through a multi-party computation protocol (henceforth called the MPC-assisted model), then we can design a finite-block TFM that simultaneously satisfies UIC, MIC, and 1-SCP (in the ex-post sense). On the other hand, Shi et al. also showed that simultaneously achieving UIC, MIC, and $c$-SCP for $c\geq 2$ is impossible in the MPC-assisted model, even for Bayesian notions of equilibrium. Partly this is because even with an MPC protocol enforcing the inclusion rule, a strategic miner or user can still inject fake bids, and the mechanism does not know the number of bids ahead of time (i.e., the mechanism is “permissionless”). It is an interesting open question whether using cryptography and Bayesian notions of equilibrium can help us overcome the impossibilities in this paper. Specifically, Shi et al. [SCW23] suggested that the MPC-assisted model also justifies the relaxed notion of Bayesian equilibrium, since players cannot see others’ bids before posting their own.
8 Static Revelation Principle for Transaction Fee Mechanisms
Informally speaking, the revelation principle says that any mechanism can be simulated by an equivalent truthful mechanism. Whenever the revelation principle holds, without loss of generality, we may assume that the users’ honest bidding strategy is simply truth-telling. We focus here on a revelation principle for “static” mechanisms in which users interact with a mechanism simply by submitting a bid. [^10] More generally, typical revelation principles in auction theory transform non-direct mechanisms (in which the action space may be different from the valuation space, for example due to multiple rounds of user-mechanism interaction) into a direct mechanisms.
TFMs differ from traditional auctions in that they additionally need to satisfy MIC and collusion-resilience, and also the separation of the inclusion rule (executed by the miner) and the confirmation/payment rules (executed by the blockchain). Therefore, revelation principles are not “automatic.” In this section, we prove that for any fixed $c$, the revelation principle holds for TFMs that must satisfy UIC, MIC, and $c$-SCP. As a direct corollary, the revelation principle holds for TFMs that satisfy UIC, MIC, and global SCP. The main subtlety in the proof is that when we bake the user’s non-truth-telling bidding rule $\beta$ into the mechanism itself, not only do we need to have the miner’s inclusion rule execute $\beta$, the blockchain’s confirmation/payment rules must also double-check the correct enforcement of $\beta$ again. Otherwise, a strategic miner may not honestly enforce $\beta$.
Notice that we focus on the mechanisms that satisfy $c$-SCP instead of OCA-proofness in this section. $c$-SCP requires that following the honest bidding rule and the honest inclusion rule is a dominant strategy for a coalition consisting of the miner and at most $c$ users. For a non-truthful mechanism with the bidding rule $\beta$, $\beta$ is used by users to satisfy both the UIC and the $c$-SCP conditions. In Section 8.1, we assume the honest bidding rule $\beta$ only outputs a single bid. Later in Section 8.2, we generalize the proof so that $\beta$ can output any non-negative number of bids.
8.1 Static Revelation Principle: Bidding Rules That Output Single Bid
The outcome of a TFM is defined as a tuple $(\mathbf{x},{\mathbf{p}},\mu)$ where $\mathbf{x}\in\{0,1\}^{t}$ is a bit-vector indicating whether each bid is confirmed or not, ${\mathbf{p}}\in\mathbb{R}_{\geq 0}^{t}$ is the vector of payments for all bids, and $\mu\in\mathbb{R}_{\geq 0}$ is the miner’s revenue.
Theorem 8.1 (Static revelation principle).
Let $c$ be any natural number. Suppose $\Pi$ is a non-truthful TFM for block size $k$ that is UIC, MIC, and $c$-SCP with an individually rational bidding rule $\beta$. If $\beta$ always outputs a single bid, then there exists a truthful TFM $\Pi^{\prime}$ for block size $k$ that is UIC, MIC, and $c$-SCP such that 1) the honest bidding rule is truth-telling, and 2) given any vector of true values, the outcome under an honest execution of $\Pi$ is identically distributed as the outcome under an honest execution of $\Pi^{\prime}$.
Proof.
Suppose we are given a TFM $\Pi=(\beta,\mathbf{I},\mathbf{C},\mathbf{P},\mu)$ where $\beta:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}$ denotes the user’s honest bidding rule, $\mathbf{I}$ denotes the inclusion rule, $\mathbf{C}$ denotes the confirmation rule, $\mathbf{P}$ denotes the payment rule, and $\mu$ denotes the miner revenue rule. Given a vector ${\mathbf{b}}=(b_{1},\dots,b_{t})$, let $\beta({\mathbf{b}})$ denote the element-wise application of $\beta$; that is, $\beta({\mathbf{b}})=(\beta(b_{1}),\dots,\beta(b_{t}))$. We construct a truthful mechanism $\Pi^{\prime}=(\mathbf{I}^{\prime},\mathbf{C}^{\prime},\mathbf{P}^{\prime},\mu^{\prime})$ as follows:
  • Inclusion rule $\mathbf{I}^{\prime}$: Given a bid vector ${\mathbf{b}}$, the miner selects the bids by using $\mathbf{I}$ as if the bid vector is $\beta({\mathbf{b}})$. That is, the miner selects the bids from ${\mathbf{b}}$ if the corresponding bids in $\beta({\mathbf{b}})$ are selected by $\mathbf{I}$.
  • Confirmation rule $\mathbf{C}^{\prime}$, payment rule $\mathbf{P}^{\prime}$ and miner revenue rule $\mu^{\prime}$: Given a created block $\mathbf{B}$, the mechanism applies $\mathbf{C}$ and $\mathbf{P}$ to $\beta(\mathbf{B})$. Then, a bid $b$ in the original block $\mathbf{B}$ is confirmed if and only if its corresponding bid $\beta(b)$ is confirmed when applying $\mathbf{C}$ to $\beta(\mathbf{B})$, and its payment is the same as its corresponding bid when applying $\mathbf{P}$ to $\beta(\mathbf{B})$. The miner revenue is $\mu(\beta(\mathbf{B}))$; that is, applying $\mu$ as if the block is $\beta(\mathbf{B})$.
Notice that $\beta$ is individually rational under $\Pi$, so the payment corresponding to the bid $\beta(v)$ never exceeds $v$. Thus, under the induced mechanism $\Pi^{\prime}$, the payment never exceeds the bid, so the individual rationality of the payment rule is respected.
Next, by directly checking the syntax, the user’s confirmation probability and expected payment when submitting a bid $b$ under $\Pi^{\prime}$ is identically distributed to submitting a bid $\beta(b)$ under $\Pi$ assuming the miner follows the inclusion rule. Similarly, the miner’s revenue when creating a block $\mathbf{B}$ under $\Pi^{\prime}$is identically distributed to creating a block $\beta(\mathbf{B})$ under $\Pi$. Thus, the outcome under an honest execution of $\Pi$ is identically distributed as the outcome under an honest execution of $\Pi^{\prime}$.
Finally, we show that the induced mechanism $\Pi^{\prime}$ satisfies UIC, MIC, and global SCP. Since submitting a bid $b$ and creating a block $\mathbf{B}$ under $\Pi^{\prime}$ is equivalent to submitting a bid $\beta(b)$ and creating a block $\beta(\mathbf{B})$ under $\Pi$, respectively, the fact that $\Pi^{\prime}$ satisfies UIC and global SCP directly follows from the fact that $\Pi$ satisfies UIC and global SCP under the user’s honest bidding rule $\beta$.
For proving $\Pi^{\prime}$ satisfies MIC, suppose $\Pi^{\prime}$ is not MIC for the sake of contradiction. In this case, there exists a bid vector ${\mathbf{b}}$ and a block $\tilde{\mathbf{B}}$ such that if the miner creates the block $\tilde{\mathbf{B}}$ instead of $\mathbf{I}^{\prime}({\mathbf{b}})$, the miner’s utility increases. Then, consider another miner under the mechanism $\Pi$, and imagine that the bid vector is $\beta({\mathbf{b}})$. Notice that $\mu(\beta(\tilde{\mathbf{B}}))=\mu^{\prime}(\tilde{\mathbf{B}})$ and $\mu(\mathbf{I}(\beta({\mathbf{b}})))=\mu^{\prime}(\mathbf{I}^{\prime}({\mathbf{b}}))$. Thus, if the miner creates the block $\beta(\tilde{\mathbf{B}})$ instead of $\mathbf{I}(\beta({\mathbf{b}}))$, the miner’s utility increases. It violates the fact that $\Pi$ is MIC. Thus, $\Pi^{\prime}$ must satisfy MIC.
8.2 Static Revelation Principle: Allowing Bidding Rules that Output Multiple Bids
In this section, we extend Theorem 8.1 to allow the bidding rule to output multiple bids. The basic idea for proving the revelation principle is still the same — trying to bake the non-truth-telling bidding rule $\beta$ into the mechanism itself. However, when $\beta$ may output multiple bids, a user $i$ will submit a vector ${\mathbf{b}}_{i}$ of bids instead of a single real number, and the inclusion rule may ask the miner to only select a subset of the bids from ${\mathbf{b}}_{i}$. Therefore, the new challenge in this case is how to make the blockchain’s confirmation/payment rules also double-check the correct enforcement of $\beta$ again, given that the blockchain can only see a subset of the bids from ${\mathbf{b}}_{i}$.
To handle this difficulty, we need to relax the syntax of the inclusion rule. Earlier in our definitions (Section 2.1), we require that the inclusion rule outputs a subset of the input bid vector. In this section, we slightly relax this syntax requirement, and allow the honest inclusion rule to append some metadata for each bid. [^11] We assume the metadata do not occupy the block space. See Remark 3 for more discussion. The purpose of the metadata in the honest execution is to encode the auxiliary information for the blockchain’s confirmation/payment rules. This relaxation in the syntax is non-essential for the impossibility results in this paper, since all the proofs for our impossibility results (Sections 4, 5 and 6.2) still hold even if the honest inclusion rule can append the metadata. Now, we are ready to present the revelation principle.
Theorem 8.2.
Let $c$ be any natural number. Given any non-truthful TFM $\Pi$ for block size $k$ that is UIC, MIC, and $c$-SCP with an individually rational bidding rule, there exists a truthful TFM $\Pi^{\prime}$ for block size $k$ that is UIC, MIC, and $c$-SCP such that 1) the honest bidding rule is truth-telling, and 2) given any vector of true values, the outcome under an honest execution of $\Pi$ is identically distributed as the outcome under an honest execution of $\Pi^{\prime}$.
Proof of Theorem 8.2.
The rest of this section is dedicated to proving Theorem 8.2. Suppose we are given a TFM $\Pi=(\beta,\mathbf{I},\mathbf{C},\mathbf{P},\mu)$ where $\beta:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{*}$ denotes the user’s honest bidding strategy, $\mathbf{I}$ denotes the inclusion rule, $\mathbf{C}$ denotes the confirmation rule, $\mathbf{P}$ denotes the payment rule, and $\mu$ denotes the miner revenue rule. We construct a truthful mechanism $\Pi^{\prime}=(\mathbf{I}^{\prime},\mathbf{C}^{\prime},\mathbf{P}^{\prime},\mu^{\prime})$ as follows:
  • Inclusion rule $\mathbf{I}^{\prime}$: given a bid vector ${\mathbf{b}}=(b_{1},\dots,b_{t})$, let $\beta({\mathbf{b}})=\left(\beta(b_{1}),\ldots,\beta(b_{t})\right)$. Create a block ${\bf B}^{\prime}$ that includes the following:
    • for each $j\in[t]$, if some bid in the vector $\beta(b_{j})$ is selected by the original inclusion rule $\mathbf{I}(\beta({\mathbf{b}}))$, then include $b_{j}$ in $\mathbf{B}^{\prime}$;
    • for each included bid $b_{j}$, attach the following extra annotation in $\mathbf{B}^{\prime}$:
      ${\sf info}(b_{j})=\text{bids in $\beta(b_{j})$ selected by $\mathbf{I}(\beta({\mathbf{b}}))$}$
    A valid block $\mathbf{B}$ is one in which the annotation for each $b_{j}\in\mathbf{B}$ is a subset of the vector $\beta(b_{j})$.
  • The confirmation rule $\mathbf{C}^{\prime}$, the payment rule $\mathbf{P}^{\prime}$, and the miner revenue rule $\mu^{\prime}$: If the input block is not valid, then no one is confirmed, and the miner revenue is zero. Else, parse the block as $\{b^{\prime}_{j},{\bf s}^{\prime}_{j}={\sf info}(b^{\prime}_{j})\}_{j}$, form an imaginary block $\{{\bf s}^{\prime}_{j}\}_{j}$, and run the original confirmation rule $\mathbf{C}$, payment rule $\mathbf{P}$ and miner revenue rule $\mu$ on $\{{\bf s}^{\prime}_{j}\}_{j}$. A bid $b^{\prime}_{j}$ in the input block is considered confirmed if the primary bid in $\beta(b^{\prime}_{j})$ is contained in ${\bf s}^{\prime}_{j}$ and is confirmed by $\mathbf{C}(\{{\bf s}^{\prime}_{j}\}_{j})$, and its payment is the sum of all payments of ${\bf s}^{\prime}_{j}$ in $\mathbf{P}(\{{\bf s}^{\prime}_{j}\}_{j})$. The miner is paid $\mu(\{{\bf s}^{\prime}_{j}\}_{j})$.
More intuitively, in the truthful mechanism $\Pi^{\prime}$, the miner simulates $\beta$ by applying $\beta$ to all bids it has received, resulting in $\beta({\mathbf{b}})$. It then simulates the original inclusion rule of the non-truthful mechanism, and the result is $\mathbf{I}(\beta({\mathbf{b}}))$, indicating which of $\beta({\mathbf{b}})$ would have been included by the original $\mathbf{I}$. The miner includes in the block each bid $b_{j}$ such that at least one of $\beta(b_{j})$ would have been included by $\mathbf{I}(\beta({\mathbf{b}}))$, and for each included bid $b_{j}$, it attaches the extra information which bids among $\beta(b_{j})$ would have been included by $\mathbf{I}(\beta({\mathbf{b}}))$. This way, the blockchain’s confirmation, payment, and miner-revenue rules can double-check the correct enforcement of $\beta$ again, before simulating the effect of the original confirmation, payment, and miner revenue rules of the non-truthful mechanism — formally, the checking is accomplished through a block validity check in our reduction.
Lemma 8.3.
Assume the miner follows the mechanism honestly. Then, the outcome when all users follow the bidding rule $\beta$ under the mechanism $\Pi$ is identically distributed as the outcome when all users bid truthfully under the mechanism $\Pi^{\prime}$.
Proof.
Let $(v_{1},\dots,v_{t})$ be the true values of all users. Under mechanism $\Pi$, if all users follow the bidding rule $\beta$, the created block will be $\mathbf{B}=\mathbf{I}(\beta(v_{1}),\dots,\beta(v_{t}))$. For any user $i$, since the fake bids have no intrinsic values, user $i$’s bid is considered as confirmed if and only if its primary bid in $\beta(v)$ is confirmed. User $i$’s payment is the sum of the payments of all its bids $\beta(v)$. The miner revenue in this case is $\mu(\mathbf{B})$.
Under mechanism $\Pi^{\prime}$, if all users bid truthfully, the created block will be $\mathbf{B}^{\prime}$ parsed as $\{b^{\prime}_{j},{\sf info}(b^{\prime}_{j})\}_{j}$. Notice that $\{{\sf info}(b^{\prime}_{j})\}_{j}$ is identically distributed as $\mathbf{B}$. The bid $b^{\prime}_{j}$ under mechanism $\Pi^{\prime}$ is considered confirmed if the primary bid in $\beta(b^{\prime}_{j})$ is contained in ${\sf info}(b^{\prime}_{j})$ and is confirmed by $\mathbf{C}(\{{\sf info}(b^{\prime}_{j})\}_{j})$, and its payment is the sum of all payments of ${\bf s}^{\prime}_{j}$ in $\mathbf{P}(\{{\sf info}(b^{\prime}_{j})\}_{j})$. The miner revenue in this case is $\mu(\{{\sf info}(b^{\prime}_{j})\}_{j})$.
Because $\{{\sf info}(b^{\prime}_{j})\}_{j}$ is identically distributed as $\mathbf{B}$, the outcomes of the honest execution of two mechanisms are also identically distributed by checking all the syntax above.
Lemma 8.4.
The following statements hold.
  • If $\Pi$ is UIC, then $\Pi^{\prime}$ is UIC.
  • If $\Pi$ is MIC, then $\Pi^{\prime}$ is MIC.
  • For all $c$, if $\Pi$ is $c$-SCP, then $\Pi^{\prime}$ is $c$-SCP.
Proof.
We will prove three properties individually.
$\Pi^{\prime}$ is UIC.
For the sake of contradiction, suppose $\Pi^{\prime}$ is not UIC. That is, under mechanism $\Pi^{\prime}$, there exists a user $i$ with true value $v$ and a bid vector ${\mathbf{b}}_{-i}$ such that user $i$’s utility increases if it submits a vector $\mathbf{f}$ (possibly includes some fake bids) instead of $v$ when other users’ bids are ${\mathbf{b}}_{-i}$. Denote the above as Scenario 1. Then, consider another Scenario 2, where the mechanism is $\Pi$, and other users’ bids are $\beta({\mathbf{b}}_{-i})$. The strategic user $i$ can submit a vector $\beta(\mathbf{f})$ instead of $\beta(v)$.
By Lemma 8.3, the confirmation probability and the payment of the bids $\mathbf{f}$ ($v$, resp.) in Scenario 1 are identically distributed as the confirmation probability and the payment of the bids $\beta(\mathbf{f})$ ($\beta(v)$, resp.) in Scenario 2. Thus, the expected utility of user $i$ if it submits $\mathbf{f}$ ($v$, resp.) in Scenario 1 is the same as the expected utility of user $i$ if it submits $\beta(\mathbf{f})$ ($\beta(v)$, resp.) in Scenario 2. Thus, in Scenario 2, user $i$’s utility increases if it $\beta(\mathbf{f})$ instead of $\beta(v)$. It violates UIC under $\Pi$, so $\Pi^{\prime}$ must satisfy UIC.
$\Pi^{\prime}$ is MIC.
For the sake of contradiction, suppose $\Pi^{\prime}$ is not MIC. That is, there exists a bid vector ${\mathbf{b}}$ and a block $\tilde{\mathbf{B}}=\{b^{\prime}_{j},{\bf s}^{\prime}_{j}={\sf info}(b^{\prime}_{j})\}_{j}$ such that if the miner creates the block $\tilde{\mathbf{B}}$ instead of $\mathbf{I}^{\prime}({\mathbf{b}})$, the miner’s utility increases. Denote the above as Scenario 1. Then, consider another Scenario 2, where the mechanism is $\Pi$, and the incoming bid vector is $\beta({\mathbf{b}})$. The strategic miner can create a block $\{{\bf s}^{\prime}_{j}\}_{j}$ instead of $\mathbf{I}(\beta({\mathbf{b}}))$.
Recall the miner’s utility is the revenue minuses the cost of injecting the fake bids. The miner revenue of the block $\tilde{\mathbf{B}}$ ($\mathbf{I}^{\prime}({\mathbf{b}})$, resp.) in Scenario 1 is the same as the miner revenue of the block $\{{\bf s}^{\prime}_{j}\}_{j}$ ($\mathbf{I}(\beta({\mathbf{b}}))$, resp.) in Scenario 2. If $\{{\bf s}^{\prime}_{j}\}_{j}$ contains some fake bids, the expected payment of those fake bids in Scenario 2 is also the same as the expected payment of the fake bids when the created block is $\tilde{\mathbf{B}}$ in Scenario 1. Therefore, in Scenario 2, if the miner creates $\{{\bf s}^{\prime}_{j}\}_{j}$ instead of $\mathbf{I}(\beta({\mathbf{b}}))$, the miner’s utility increases. It violates MIC under $\Pi$, so $\Pi^{\prime}$ must satisfy MIC.
$\Pi^{\prime}$ is $c$-SCP for all $c$.
For the sake of contradiction, suppose $\Pi^{\prime}$ is not $c$-SCP for some $c$. That is, there exist a coalition $C$ formed by the miner and at most $c$ users, a bid vector ${\mathbf{b}}_{-C}$, and a block $\tilde{\mathbf{B}}$ such that when all non-colluding users’ bids are ${\mathbf{b}}_{-C}$, $C$’s utility increases if the created block is $\tilde{\mathbf{B}}$ instead of $\mathbf{I}^{\prime}({\mathbf{b}}_{-C},{\mathbf{b}}_{C})$, where ${\mathbf{b}}_{C}$ is the true values of all users in $C$. Denote the above as Scenario 1. Then, consider another Scenario 2, which is the same as scenario 1 except that the mechanism is $\Pi$, and the bid vector of non-colluding user is $\beta({\mathbf{b}}_{-C})$. Parse $\tilde{\mathbf{B}}=\{b^{\prime}_{j},{\bf s}^{\prime}_{j}={\sf info}(b^{\prime}_{j})\}_{j}$. The coalition $C$ can create a block $\{{\bf s}^{\prime}_{j}\}_{j}$ instead of $\mathbf{I}\left(\beta({\mathbf{b}}_{-C},{\mathbf{b}}_{C})\right)$.
Following the same argument as UIC and MIC, we conclude that the joint utility of $C$ of the block $\tilde{\mathbf{B}}$ ($\mathbf{I}^{\prime}({\mathbf{b}}_{-C},{\mathbf{b}}_{C})$, resp.) in Scenario 1 is the same as the joint utility of $C$ of the block $\{{\bf s}^{\prime}_{j}\}_{j}$ ($\mathbf{I}(\beta({\mathbf{b}}_{-C},{\mathbf{b}}_{C}))$, resp.) in Scenario 2. Thus, in Scenario 2, if the coalition creates $\{{\bf s}^{\prime}_{j}\}_{j}$ instead of $\mathbf{I}\left(\beta({\mathbf{b}}_{-C},{\mathbf{b}}_{C})\right)$, the joint utility of $C$ increases. It violates $c$-SCP under $\Pi$, so $\Pi^{\prime}$ must satisfy $c$-SCP. ∎
The proof of Theorem 8.2 directly follows from Lemma 8.3 and Lemma 8.4.
Remark 3.
We assume the metadata appended by the inclusion rule do not occupy the block space. If the metadata occupy the block space, it is not hard to see that we can encode a valid block using a vector of length at most $3k$. For example, we can divide the block into three parts: the first part encodes users’ original bids $b_{j}$; the second part encodes $\beta(b_{j})$; and the third part encodes the mapping between the first part and the second part.
Appendix A Comparison of Collusion-Resilience Notions
In this work, we discuss three different definitions of collusion-resilience: $c$-SCP, global SCP, and OCA-proofness. Both global SCP and OCA-proofness focus on the global coalition consisting of the miner and all users, while $c$-SCP considers a coalition formed by the miner and at most $c$ users. Both $c$-SCP and global SCP require following the honest bidding rule is a dominant strategy, so if a TFM satisfies $c$-SCP for all $c$, it must also satisfy global SCP. On the other hand, an OCA-proof TFM allows each user to bid according to an individually rational bidding strategy $\sigma:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}$ as long as $\sigma$ does not depend on others’ bids. Hence, for any TFM whose bidding rule only outputs a single real, the TFM satisfies global SCP implies it also satisfies OCA-proofness by simply taking the honest bidding rule as the individually rational bidding strategy $\sigma$. However, in Section 2.1, the honest bidding rule can output multiple bids in general. In this case, global SCP may not imply OCA-proofness. Consider the following somewhat contrived mechanism.
{mdframed}
Even Auction
Parameters: the block size is infinite.
Mechanism:
  • Bidding rule. Each user submits two bids $(v,0)$, where $v$ is the true value.
  • Inclusion rule. Include all pending bids.
  • Confirmation rule. If the number of the included bids is odd, no one is confirmed. Otherwise, if the number of the included bids is even, all included bids are confirmed.
  • Payment rule and miner revenue rule. Everyone pays nothing, and the miner revenue is zero.
Lemma A.1.
The even auction above satisfies global SCP but not OCA-proofness.
Proof.
If everyone follows the bidding rule, the number of the included bids must be even, so everyone’s bid is confirmed. Thus, the social welfare is maximized, and the mechanism satisfies global SCP. However, when there are an odd number of users, following any bidding strategy $\sigma:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}$ will lead to an odd number of bids, and no one is confirmed. Thus, it does not satisfy OCA-proofness. ∎
Notice that whether the individually rational bidding strategy $\sigma$ outputs a single bid or multiple bids is important. If we require $\sigma$ to only output a single real as we defined in Definition 6, there is no non-trivial TFM that can satisfy UIC, MIC, and OCA-proofness as we proved in Theorem 6.9. However, if we allow $\sigma$ to output multiple bids, we obtain a feasibility in Section 7.2.
When we analyze whether a TFM satisfies $c$-SCP for some $c$, $1$-SCP is the weakest requirement since $c^{\prime}$-SCP always implies $c$-SCP for any $c^{\prime}>c$ by definition. Interestingly, even so, $1$-SCP is incomparable with global SCP and OCA-proofness. The relations between $1$-SCP, global SCP, and OCA-proofness is depicted in Figure 1.
[…]
Figure 1: Relationship between collusion-resilience notions.
We explain Figure 1 in more detail below:
  • global SCP $\overset{\triangle}{\implies}$ OCA-proofness: As we discussed above, if the honest bidding rule only outputs a single real, global SCP implies OCA-proofness; if the honest bidding rule outputs multiple reals, Lemma A.1 states that global SCP does not imply OCA-proofness.
[…]
Below, we introduce the (somewhat contrived) auctions for constructing the counterexamples.
{mdframed}
Pay-Nothing Auction
Parameters: the block size $k$.
Mechanism:
  • Bidding rule. Each user submits the true value.
  • Inclusion rule. Choose the highest up to $k$ bids.
  • Confirmation rule, payment rule and miner revenue rule. All included bids are confirmed, everyone pays nothing, and the miner revenue is zero.
Lemma A.2.
The pay-nothing auction above satisfies global SCP and OCA-proofness. However, it does not satisfy $1$-SCP.
Proof.
When everyone follows the mechanism, the users with the highest values will be confirmed, so the social welfare is maximized. Thus, the mechanism satisfies global SCP and OCA-proofness. However, if there are $k+1$ users with the true values $v_{1}>\cdots>v_{k+1}>0$. When everyone follows the mechanism, the $(k+1)$-th user’s bid will not be included in the block, so the user’s utility is zero. The platform can collude with the $(k+1)$-th user, and include its bid. Since the payment is always zero, the coalition now has positive utility, which violates $1$-SCP. ∎
{mdframed}
Discount Auction
Parameters: the fixed price $r$, and the block size is infinite.
Mechanism:
  • Bidding rule. Each user submits the true value.
  • Inclusion rule. Include all pending bids.
  • Confirmation rule. Let $S$ be the set of all included bids, and define $f$ as
    $f(t)=\left\{\begin{matrix}r,&\text{if }t\leq 10;\\ r/2,&\text{otherwise}.\end{matrix}\right.$
    Let $t^{*}$ be the largest integer such that there are $t^{*}$ bids in $S$ larger than $f(t^{*})$. All bids larger than $f(t^{*})$ are confirmed.
  • Payment rule. Each confirmed bid pays $f(t^{*})$. Unconfirmed bids pay nothing.
  • Miner revenue rule. The miner revenue is zero, and all payments are burnt.
Lemma A.3.
The discount auction above satisfies $1$-SCP. However, it does not satisfy global SCP or OCA-proofness.
Proof.
We first show that the mechanism satisfies $1$-SCP. Consider any coalition $C$ formed by the miner and user $i$. Because the miner revenue is always zero, $C$’s joint utility can only come from user $i$. Fix any bid vector ${\mathbf{b}}_{-i}$ from other users, user $i$’s confirmation and the payment satisfy the requirements in Myerson’s lemma. Thus, overbidding and underbidding cannot increase user $i$’s utility. Next, we show that injecting fake bids does not help either. To increase user $i$’s utility, the only way is to lower the payment. However, regardless of the bids, the payment for each bid is always in $[r/2,r]$, so injecting fake bids can only lower the payment by at most $r/2$. On the other hand, all confirmed bids must pay at least $r/2$, which means user $i$ has to pay at least $r/2$ for each fake bid. Thus, injecting fake bids do not increase user $i$’s utility, so we conclude that the mechanism satisfies $1$-SCP.
Now, we show that the mechanism does not satisfy OCA-proofness. Imagine a scenario where there are $10$ users each with the true value $r$. Without injecting any fake bid, the lowest possible payment for each user is $f(10)=r$, so the social welfare is at most 00. However, if the global coalition inject one fake bid $r$, everyone’s payment will further lower to $f(11)=r/2$. In this case, the social welfare becomes $010r-11r/2=9r/2>$ 0. Thus, the mechanism does not satisfy OCA-proofness.
The mechanism does not satisfy global SCP either since the bidding rule only outputs a single real. If the mechanism satisfied global SCP, the global coalition could pick the suggested bidding rule as the individually rational bidding strategy $\sigma$ to maximize the social welfare. Given that the mechanism does not satisfy OCA-proofness, it cannot satisfy global SCP either. ∎